### v2

# Null hypersurfaces in general relativity: Intrinsic symmetries and differential invariants

Abstract
This paper investigates intrinsic Killing symmetries of null hypersurfaces $\mathcal{N}_{3}$ within the framework of general relativity. To this end we consider $\mathcal{N}_{3}$ as detached from the embedding spacetime and equipped with a degenerate metric of signature (0,+,+). As geometrical tools we use a triad calculus and differential invariants. Based on prior work, we present a classification of null hypersurfaces according to groups of motion up to the fourth order. For each type certain normal forms of the metric are given, and their invariants are listed. A discussion of horizons - defined as null hypersurfaces with vanishing shear and divergence - is included.
1 Introduction
Null hypersurfaces, also referred to as lightlike, characteristic or isotropic hypersurfaces, play a central role in general relativity. Prominent examples include the formation and structure of horizons in black hole physics, the theory of gravitational radiation with characteristic propagation fronts, and the cosmological past light cone, which encodes the accessible empirical information of the Universe. Another significant example is the characteristic initial value problem of Einstein’s field equations, based on a pair of null hypersurfaces, which was first addressed many years ago [27, 24, 10]. Since then, many further approaches for studying null hypersurfaces and their embedding in spacetime have been developed, see the extensive reference lists in recent publications [9, 25, 2, 21, 22, 7, 1, 20].
Null hypersurfaces are embedded within the spacetime geometry, and geometrical quantities associated with this embedding - such as surface gravity and rigged null directions - play a crucial role in various applications. However, a clear geometric understanding can only be achieved by carefully separating intrinsic properties from embedding characteristics. Thus the null hypersurface can initially be considered as an independent object, detached from the ambient spacetime (see also [22]).
We adopt this strategy and investigate in this paper the Killing isometries of a three-dimensional standalone null hypersurface $\mathcal{N}_{3}$. A related study was carried out previously [13], but it contained many misprints and errors. Here we correct that analysis.
The article is organized as follows. In section 2 an elementary presentation of the geometry of null surfaces is given in terms of a triad formalism. A detailed derivation of differential invariants follows in section 3. Section 4 treats the Killing equation and its integrability conditions in the triad calculus. Section 5 focused on hypersurfaces admitting a null Killing vector (Killing horizons). Section 6 determines metrics admitting groups of motion $G_{1}$ and $G_{2}$. Sections 7 and 8 treat the same problem for the groups $G_{3}$ and $G_{4}$ respectively. In section 9 some properties of the metrics are briefly outlined, such as orbits of Killing vectors. Appendix A summarizes the classification of null hypersurfaces in terms of differential invariants and provides tools for most of the calculations in the article. Appendix B contains an overview of the found metrics.
The considerations in this article refer to the inner geometry, which is based solely on a degenerate metric $\gamma_{ik}$. In [28] and [11, 12] it was shown, that a whole hierarchy of null surface geometries exist. They range from the conformal geometry to specialized inner geometries, if less or more properties of the ambient spacetime are added. Another paper may address these issues.
More mathematically oriented treatments of degenerate metrics with isometries can be found in numerous publications, for example [4],[5],[26].
2 A triad formalism for $\mathcal{N}_{3}$
We use a triad calculus to study null spaces $\mathcal{N}_{3}$. $\mathcal{N}_{3}$ is defined as three-dimensional manifold with a smooth symmetric covariant tensor field $\gamma_{ik}(x_{i})$ of matrix rank 2 and signature $(0,1,1)$ as metric. Regular points $x_{i}$ are defined by $rank~\gamma_{ik}(x_{i})=2.$ The rank may drop to 1 at focal points and 0 at a vertex, but such cases will not be considered. Also global aspects of $\mathcal{N}_{3}$ are not discussed here. We mention only that the special role of null directions suggests a product topology, e.g. $R^{2}\times R$ or $S^{2}\times R$. Since $\gamma_{ik}$ is degenerate, the equation
$\gamma_{ik}\epsilon^{k}=0$ (2.1)
has a non-vanishing solution $\epsilon^{k}$, defined up to a factor. $\epsilon^{k}$ is called null direction or generator direction. The tangential lines of $\epsilon^{k}$ provide a foliation of $\mathcal{N}_{3}$. In addition to $\epsilon^{k}$ there are two further vectors $p^{k},~q^{k}$, such that
$p^{k}\gamma_{kl}=p_{l}\neq 0,~q^{k}\gamma_{kl}=q_{l}\neq 0.$ (2.2)
We combine $p^{k}$ and $q^{k}$ into a complex pseudo null vector:
$t^{k}=\frac{1}{\sqrt{2}}(p^{k}+iq^{k})$ (2.3)
The corresponding complex-conjugate vector is denoted by $\bar{t^{k}}=\frac{1}{\sqrt{2}}(p^{k}-iq^{k})$. Together with the contravariant triad $\epsilon^{k},t^{k},\bar{t^{k}}$ we introduce a corresponding covariant triad $\gamma_{k},t_{k},\bar{t_{k}}$ by
$t^{k}t_{k}=0,~t^{k}\bar{t_{k}}=1,~\epsilon^{k}t_{k}=t^{k}\gamma_{k}=0,~\epsilon^{k}\gamma_{k}=1.$ (2.4)
The metric can now be written
$\gamma_{ik}=t_{i}\bar{t_{k}}+\bar{t_{i}}t_{k}.$ (2.5)
Since $\gamma_{ik}$ is degenerate, there is no reciprocal contravariant metric $\epsilon^{ik}$ with $\epsilon^{ik}\gamma_{kl}=\delta^{i}_{l}$. For a non-degenerate metric the latter condition is equivalent to
$\gamma_{li}\gamma_{mk}\epsilon^{ik}=\gamma_{lm}.$
In the case of degeneration, this equation remains meaningful, but the solution for $\epsilon^{ik}$ is not unique, since
$\epsilon^{\prime ik}=\epsilon^{ik}+a^{i}\epsilon^{k}+a^{k}\epsilon^{i}$
is also a solution for every $a^{i}$. Nevertheless, with
$\epsilon^{ik}=\bar{t^{i}}t^{k}+\bar{t^{k}}t^{i}$ (2.6)
a contravariant metric $\epsilon^{ik}$ can be unambiguously assigned to each triad. $\epsilon^{ik}$ has the same rank as $\gamma_{ik}$ and is related to the covariant metric by
$\epsilon^{kl}\gamma_{il}=\delta^{k}_{i}-\epsilon^{k}\gamma_{i}.$ (2.7)
The triad is not unique. It can be transformed into
$t^{\prime i}=e^{i\omega}(t^{i}-\lambda\bar{\kappa}\epsilon^{i}),$ (2.8a)
$\epsilon^{\prime i}=\lambda\epsilon^{i},$ (2.8b)
$t^{\prime}_{i}=e^{i\omega}t_{i},$ (2.8c)
$\gamma^{\prime}_{i}=\frac{1}{\lambda}\gamma_{i}+\kappa t_{i}+\bar{\kappa}\bar{t}_{i},$ (2.8d)
$\omega,\lambda$ real, $\kappa$ complex. This transformation group is locally isomorphic to a 4-parameter subgroup of the Lorentz transformations that leave the null direction invariant. It splits into null rotations ($\kappa$), spacelike rotations ($\omega$) and a scale change ($\lambda$).
The introduction of covariant derivatives requires the definition of an affinity. This is well known to encounter difficulties in degenerate spaces due to the non-existence of a contravariant metric [19, 17]. We solve this problem by introducing a class of affinities, that is a triad-depending affinity (see [11],[12]). For a given triad we define the affinity by
$\Gamma^{l}_{ik}=\frac{1}{2}\epsilon^{lm}(\gamma_{mi,k}+\gamma_{mk,i}-\gamma_{ik,m})+\frac{1}{2}\epsilon^{l}(\gamma_{i,k}+\gamma_{k,i}).$ (2.8i)
Under general coordinate transformation
$x^{\prime}_{i}=x^{\prime}_{i}(x_{k})$ (2.8j)
with $det|\frac{\partial x^{\prime}_{i}}{\partial x_{k}}|\neq 0$, $\Gamma^{i}_{kl}$ behaves as affinity indeed. Using this affinity, the covariant derivatives of the triad (denoted by a stroke, for the ordinary derivative we use a comma) may be represented as:
$\epsilon^{i}_{|k}=-(\rho\bar{t}_{k}+\sigma t_{k})t^{i}-(\rho t_{k}+\bar{\sigma}\bar{t}_{k})\bar{t}^{i}+\chi t_{k}\epsilon^{i}+\bar{\chi}\bar{t}_{k}\epsilon^{i},$ (2.8k)
$t^{i}_{|k}=(\tau t_{k}-\bar{\tau}\bar{t}_{k}+i\nu\gamma_{k})t^{i}+(i\varphi t_{k}-\bar{\chi}\gamma_{k})\epsilon^{i},$ (2.8l)
$\gamma_{i|k}=i\varphi(t_{i}\bar{t}_{k}-\bar{t}_{i}t_{k})+\chi(t_{i}\gamma_{k}-\gamma_{i}t_{k})+\bar{\chi}(\bar{t}_{i}\gamma_{k}-\gamma_{i}\bar{t}_{k})$ (2.8m)
$t_{i|k}=\tau t_{i}t_{k}-\bar{\tau}t_{i}\bar{t}_{k}+i\nu\gamma_{k}t_{i}+\rho\gamma_{i}t_{k}+\bar{\sigma}\bar{t}_{k}\gamma_{i}.$ (2.8n)
The so-called rotation coefficients $\rho~,\sigma~,\nu~,\tau~,\chi~,\varphi$ can also be displayed without affinity:
$\rho+i\nu=\epsilon^{i}t^{k}(\bar{t}_{i|k}-\bar{t}_{k|i})=\epsilon^{i}t^{k}(\bar{t}_{i,k}-\bar{t}_{k,i}),$ (2.8o)
$\sigma=\epsilon^{i}\bar{t}^{k}\bar{t}_{i|k}=\epsilon^{i}\bar{t}^{k}(\bar{t}_{i,k}-\bar{t}_{k,i}),$ (2.8p)
$\tau=\bar{t}^{i}\bar{t}^{k}t_{i|k}=\bar{t}^{i}t^{k}(\bar{t}_{i,k}-\bar{t}_{k,i}),$ (2.8q)
$\chi=\bar{t}^{i}\epsilon^{k}\gamma_{i|k}=\frac{1}{2}\bar{t}^{i}\epsilon^{k}(\gamma_{i,k}-\gamma_{k,i}),$ (2.8r)
$i\varphi=\bar{t}^{i}t^{k}\gamma_{i|k}=\frac{1}{2}\bar{t}^{i}t^{k}(\gamma_{i,k}-\gamma_{k,i}).$ (2.8s)
It is appropriate to introduce directional differential operators:
$\delta=t^{i}\partial_{i},~\bar{\delta}=\bar{t}^{i}\partial_{i},~D=\epsilon^{i}\partial_{i},~\partial_{i}\equiv\frac{\partial}{\partial x_{i}}.$ (2.8t)
They satisfy the commutator relations
$D\delta-\delta D=(\rho+i\nu)\delta+\bar{\sigma}\bar{\delta}-2\bar{\chi}D,$ (2.8u)
$D\bar{\delta}-\bar{\delta}D=(\rho-i\nu)\bar{\delta}+\sigma\delta-2\chi D,$ (2.8v)
$\delta\bar{\delta}-\bar{\delta}\delta=\bar{\tau}\bar{\delta}-\tau\delta-2i\varphi D.$ (2.8w)
We also note the identities
$D\tau$ $=$ $\bar{\delta}(\rho+i\upsilon)-\delta\sigma+\tau(\rho-i\upsilon)+\sigma\bar{\tau}-2\chi(\rho+i\upsilon)+2\sigma\bar{\chi},$ (2.8x)
$iD\varphi$ $=$ $\delta\chi-\bar{\delta}\bar{\chi}-\chi\bar{\tau}+\bar{\chi}\tau+2i\varphi\rho.$ (2.8y)
Under the triad transformations (2.8a)-(2.8d) the rotation coefficients transform as
$\rho^{\prime}$ $=$ $\lambda\rho,$ (2.8z)
$\sigma^{\prime}$ $=$ $\lambda e^{-2i\omega}\sigma,$ (2.8aa)
$\tau^{\prime}$ $=$ $e^{-i\omega}(\tau+i\bar{\delta}\omega-i\kappa\lambda D\omega-i\kappa\lambda\upsilon+\bar{\kappa}\lambda\sigma-\kappa\lambda\rho),$ (2.8ab)
$\upsilon^{\prime}$ $=$ $\lambda\upsilon+\lambda D\omega,$ (2.8ac)
$\chi^{\prime}$ $=$ $\frac{1}{2}e^{-i\omega}(2\chi+\frac{\bar{\delta}\lambda}{\lambda}-\kappa\rho\lambda+i\upsilon\kappa\lambda-\bar{\kappa}\lambda\sigma+\lambda D\kappa),$ (2.8ad)
$i\varphi^{\prime}$ $=$ $\frac{1}{\lambda}i\varphi+\frac{1}{2}(\bar{\kappa}\tau-\kappa\bar{\tau}+\delta\kappa-\bar{\delta}\bar{\kappa})+\frac{1}{2}\kappa(2\bar{\chi}+\frac{\delta\lambda}{\lambda}-i\bar{\kappa}\lambda\upsilon-\kappa\lambda\bar{\sigma}+\lambda D\bar{\kappa})$ (2.8ae)
$-\frac{1}{2}\bar{\kappa}(2\chi+\frac{\bar{\delta}\lambda}{\lambda}+i\kappa\lambda\upsilon-\bar{\kappa}\lambda\sigma+\lambda D\kappa)$
We may use the transformations to simplify many calculations. Thus $\gamma_{i}$ can be chosen as gradient:
$\gamma_{i}=v_{,i.}.$ (2.8af)
In this case
$\chi=\varphi=0.$ (2.8ag)
$\mathcal{N}_{3}$ is then covered by a set of spacelike 2-slices $v=const$. The slices have a two-dimensional metric with a Gaussian curvature
$K=-2\tau\bar{\tau}+\bar{\delta}\bar{\tau}+\delta\tau.$ (2.8ah)
The condition (2.8ag) is preserved under triad transformations with coefficients $\lambda$ and $\kappa$ restricted by
$\delta\lambda=\lambda^{2}(-D\bar{\kappa}+i\bar{\kappa}\nu+\bar{\kappa}\rho+\kappa\bar{\sigma}),$ (2.8ai)
$\bar{\delta}\bar{\kappa}-\delta\kappa=\bar{\kappa}\tau-\kappa\bar{\tau}.$ (2.8aj)
In general, a triad transformation leads to a different set of space-like slices with a changed Gaussian curvature, thus $K$ is in general no differential invariant of $\mathcal{N}_{3}$ [11, 12]. An important exception are horizons, cf. Section 4.
Another quantity which can be made to disappear in regular points is $\upsilon=\ \frac{1}{i}t_{i\mid k}\epsilon^{k}\bar{t}^{i}$, which gives the deviation of the displacement of the components $t^{i}$ in the null direction from parallel transport with respect to the affinity $\Gamma^{l}_{~ik}$. $\upsilon=0$ is equivalent to parallel displacement of both the covariant and the contravariant triad along $\epsilon^{i}$. To maintain $\upsilon=0$, the triad transformations would have to be restricted by the further requirement
$D\omega=0.$ (2.8ak)
In this paper we use mainly the simplified form of the triad formalism with $\chi=\varphi=\upsilon=0.$
It is also useful to introduce adapted coordinates defined by $\epsilon^{i}=\delta^{i}_{1}$, using the transformations (2.8j) [^1] One example where all components of $\gamma_{ik}$ are required, is the null quasispherical gauge by Bartnik [3]. . Then, from (2.1), $\gamma_{i1}=0$, so that the metric can be written in the standard form
$ds^{2}=E(x_{1},x_{2},x_{3})dx_{2}^{2}+2F(x_{1},x_{2},x_{3})dx_{2}dx_{3}+G(x_{1},x_{2},x_{3})dx_{3}^{2}$ (2.8al)
with three functions $E,F,G.$ Adapted coordinates remain adapted if the transformations (2.8j) are restricted to (capital letters always run from 2 to 3)
$x^{\prime}_{1}=x^{\prime}_{1}(x_{1},x_{A}),$ (2.8am)
$x^{\prime}_{A}=x^{\prime}_{A}(x_{B}).$ (2.8an)
3 Differential invariants
Differential invariants are an important geometric element of null spaces. The invariants are functions of the metric and its derivatives up to a certain order:
[…]
that are invariant under the transformation (2.8j). We can also represent the invariants by rotation coefficients and their directional derivatives. In this paper we consider only invariants up to second order and write therefore
$f=f(\rho,\sigma,\tau,\nu,\chi,\varphi,D\rho,...\bar{\delta}\varphi).$ (2.8b)
The function $f$ must be invariant under the triad transformations (2.8a)-(2.8d). We use the restricted set of rotation coefficients and will first consider first-order invariants depending on $\rho,\sigma,\bar{\sigma},\tau,\bar{\tau}$:
$f(\rho,\sigma,\bar{\sigma},\tau,\bar{\tau})=f(\rho^{\prime},\sigma^{\prime},\bar{\sigma^{\prime}},\tau^{\prime},\bar{\tau^{\prime}}).$ (2.8c)
Using infinitesimal transformations
$\lambda=1+\check{\lambda},~e^{i\omega}=1+i\check{\omega},~\kappa=\check{\kappa}$ (2.8d)
with small $\check{\lambda},\check{\omega},\check{\kappa}$ and expanding the rhs to first order, we obtain:
$\check{\lambda}(\rho\frac{\partial f}{\delta\rho}+\sigma\frac{\delta f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}})+i\check{\omega}(-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}})$
$+\check{\kappa}(-\rho\frac{\partial f}{\partial\tau}+\bar{\sigma}\frac{\partial f}{\partial\bar{\tau}})+\check{\bar{\kappa}}(-\rho\frac{\partial f}{\partial\bar{\tau}}+\sigma\frac{\partial f}{\partial\tau})+i\bar{\delta}\check{\omega}\frac{\partial f}{\partial\tau}-i\delta\check{\omega}\frac{\partial f}{\partial\bar{\tau}}=0.$
The relations must be satisfied for all $\check{\lambda},\check{\omega},\check{\kappa}$, leading to
$\frac{\partial f}{\partial\tau}=0,~\frac{\partial f}{\partial\bar{\tau}}=0,~\rho\frac{\partial f}{\partial\rho}+\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}=0,~\sigma\frac{\partial f}{\partial\sigma}-\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}=0.$ (2.8e)
These are essentially $m=2$ linear partial differential equations for $n=3$ independent variables $\rho,\sigma,\bar{\sigma}$. Systems of differential equations of this type are treated, e.g., by Kamke [18]. They are called involution systems, if the integrability conditions (IC) are satisfied. With the derivative operators
$F^{1}=\rho\frac{\partial}{\partial\rho}+\sigma\frac{\partial}{\partial\sigma}+\bar{\sigma}\frac{\partial}{\partial\bar{\sigma}},~F^{2}=\sigma\frac{\partial}{\partial\sigma}-\bar{\sigma}\frac{\partial}{\partial\bar{\sigma}}$ (2.8f)
we write differential equations for $f$
$F^{i}f=0,i=1...m.$ (2.8g)
The m(m1)/2=1m(m-1)/2=1 IC’s are then
$F^{i}(F^{k})f-F^{k}(F^{i})f=0.$ (2.8h)
In our case the IC’s are satisfied. When considering the number of solutions, the rank of the coefficient matrix plays a role. This is the $m\times n$ matrix:
$\left[{\begin{array}[]{ccc}\rho&\sigma&\bar{\sigma}\\ 0&\sigma&-\bar{\sigma}\end{array}}\right]$ (2.8i)
The rank of this matrix is 2, if $|\sigma|\neq 0$. In this case nm=1n-m=1 function forms an integral basis. Obviously
$j=\rho/|\sigma|$ (2.8j)
or any function of j represent the first order invariant. Only this first-order invariant exists.
Turning now to second-order invariants, we can write
$f(D\rho,D\sigma,D\bar{\sigma},\delta\rho,\bar{\delta}\rho,\delta\sigma,\bar{\delta}\sigma,\delta\bar{\sigma},\bar{\delta}\bar{\sigma},\delta\tau,\bar{\delta}\tau,\delta\bar{\tau},\bar{\delta}\bar{\tau},\rho,\sigma,\bar{\sigma},\tau,\bar{\tau})=$
$f(D^{\prime}\rho^{\prime},D^{\prime}\sigma^{\prime},D^{\prime}\bar{\sigma^{\prime}},\delta^{\prime}\rho^{\prime},\bar{\delta^{\prime}}\rho^{\prime},\delta^{\prime}\sigma^{\prime},\bar{\delta^{\prime}}\sigma^{\prime},\delta^{\prime}\bar{\sigma^{\prime}},\bar{\delta^{\prime}}\bar{\sigma^{\prime}},\delta^{\prime}\tau^{\prime},\bar{\delta^{\prime}}\tau^{\prime},\delta^{\prime}\bar{\tau^{\prime}},\bar{\delta^{\prime}}\bar{\tau^{\prime}},$ (2.8k)
$\rho^{\prime},\sigma^{\prime},\bar{\sigma^{\prime}},\tau^{\prime},\bar{\tau^{\prime}}).$
Because of the identity (2.8x), $f$ may be considered as independent of $D\tau,D\bar{\tau}$, also the simplified version of the triad is taken. With the infinitesimal transformations (2.8d) we have to linear order
$D^{\prime}\rho^{\prime}=D\rho+\rho D\check{\lambda}+2\check{\lambda}D\rho,$ (2.8l)
$D^{\prime}\sigma^{\prime}=D\sigma+\sigma D\check{\lambda}-2i\check{\omega}D\sigma+2\check{\lambda}D\sigma,$ (2.8m)
$\delta^{\prime}\rho^{\prime}=\delta\rho+\rho\delta\check{\lambda}+\check{\lambda}\delta\rho+i\check{\omega}\delta\rho-\check{\bar{\kappa}}D\rho,$ (2.8n)
$\delta^{\prime}\sigma^{\prime}=\delta\sigma+\check{\lambda}\delta\sigma+\sigma\delta\check{\lambda}-2i\sigma\delta\check{\omega}-i\check{\omega}\delta\sigma-\check{\bar{\kappa}}D\sigma,$ (2.8o)
$\bar{\delta}^{\prime}\sigma^{\prime}=\bar{\delta}\sigma+\check{\lambda}\bar{\delta}\sigma+\sigma\bar{\delta}\check{\lambda}-2i\sigma\bar{\delta}\check{\omega}-3i\check{\omega}\bar{\delta}\sigma-\check{\kappa}D\sigma,$ (2.8p)
$\delta^{\prime}\tau^{\prime}=\delta\tau-i\tau\delta\check{\omega}+i\delta\bar{\delta}\check{\omega}+\sigma\delta\check{\bar{\kappa}}+\check{\bar{\kappa}}\delta\sigma-\rho\delta\check{\kappa}-\check{\kappa}\delta\rho-\check{\bar{\kappa}}D\tau,$ (2.8q)
$\bar{\delta}^{\prime}\tau^{\prime}=\bar{\delta}\tau-2i\check{\omega}\bar{\delta}\tau-i\tau\bar{\delta}\check{\omega}+i\bar{\delta}\bar{\delta}\check{\omega}+\sigma\bar{\delta}\check{\bar{\kappa}}+\check{\bar{\kappa}}\bar{\delta}\sigma-\rho\bar{\delta}\check{\kappa}-\check{\kappa}\bar{\delta}\rho-\check{\kappa}D\tau,$ (2.8r)
$\rho^{\prime}=\rho+\check{\lambda}\rho,$ (2.8s)
$\sigma^{\prime}=\sigma+\check{\lambda}\sigma-2i\check{\omega}\sigma,$ (2.8t)
$\tau^{\prime}=\tau-i\check{\omega}\tau+i\bar{\delta}\check{\omega}+\sigma\check{\bar{\kappa}}-\rho\check{\kappa}.$ (2.8u)
Expanding 3, we get
$\check{\lambda}\Bigl(2D\rho\frac{\partial f}{\partial D\rho}+2D\sigma\frac{\partial f}{\partial D\sigma}+2D\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}+\delta\rho\frac{\partial f}{\partial\delta\rho}+\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\rho}+\delta\sigma\frac{\partial f}{\partial\delta\sigma}$
$+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+\rho\frac{\partial f}{\partial\rho}+\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}\Bigr)$
$+D\check{\lambda}\Bigl(\rho\frac{\partial f}{\partial D\rho}+\sigma\frac{\partial f}{\partial D\sigma}+\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}\Bigr)$
$+i\check{\omega}\Bigl(-2D\sigma\frac{\partial f}{\partial D\sigma}+2D\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}+\delta\rho\frac{\partial f}{\partial\delta\rho}-\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\rho}-\delta\sigma\frac{\partial f}{\partial\delta\sigma}$
$-3\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+3\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+2\delta\bar{\tau}\frac{\partial f}{\partial\delta{\bar{\tau}}}-2\bar{\delta}\tau\frac{\partial f}{\partial\bar{\delta}\tau}$
$-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}\Bigr)$
$+i\delta\check{\omega}\Bigl(-\tau\frac{\partial f}{\partial\delta\tau}+\bar{\tau}\frac{\partial f}{\partial\delta\bar{\tau}}-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}\Bigr)$
$+i\bar{\delta}\check{\omega}\Bigl(-\tau\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}+\frac{\partial f}{\partial\tau}-2\sigma\frac{\partial f}{\partial\bar{\delta\sigma}}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}\Bigr)$
$+\delta\check{\lambda}\Bigl(\rho\frac{\partial f}{\partial\delta\rho}+\sigma\frac{\partial f}{\partial\delta\sigma}+\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}\Bigr)+\bar{\delta}\check{\lambda}\Bigl(\rho\frac{\partial f}{\partial\bar{\delta}\rho}+\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}\Bigr)$
$+\check{\bar{\kappa}}\Bigl(-D\rho\frac{\partial f}{\partial\delta\rho}-D\sigma\frac{\partial f}{\partial\delta\sigma}-D\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\delta\sigma\frac{\partial f}{\partial\delta\tau}-D\tau\frac{\partial f}{\partial\delta\tau}-\delta\rho\frac{\partial f}{\partial\delta\bar{\tau}}$
$-D\bar{\tau}\frac{\partial f}{\partial\delta\bar{\tau}}+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\tau}-\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}+\sigma\frac{\partial f}{\partial\tau}-\rho\frac{\partial f}{\partial\bar{\tau}}\Bigr)$
$+\check{\kappa}\Bigl(-D\rho\frac{\partial f}{\partial\bar{\delta}\rho}-D\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}-D\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-\delta\rho\frac{\partial f}{\partial\delta\tau}-D\bar{\tau}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\tau}}$
$-D\tau\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}-\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\tau}-\rho\frac{\partial f}{\partial\tau}+\bar{\sigma}\frac{\partial f}{\partial\bar{\tau}}\Bigr)$
$+\delta\check{\bar{\kappa}}\Bigl(\sigma\frac{\partial f}{\partial\delta\tau}-\rho\frac{\partial f}{\partial\delta\bar{\tau}}\Bigr)+\delta\check{\kappa}\Bigl(\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\tau}}-\rho\frac{\partial f}{\partial\delta\tau}\Bigr)$
$+\bar{\delta}\check{\kappa}\Bigl(-\rho\frac{\partial f}{\partial\bar{\delta}\tau}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}\Bigr)+\bar{\delta}\check{\bar{\kappa}}\Bigl(\sigma\frac{\partial f}{\partial\bar{\delta}\tau}-\rho\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}\Bigr)$
$-i\delta\delta\check{\omega}\frac{\partial f}{\partial\delta\bar{\tau}}+i\bar{\delta}\bar{\delta}\check{\omega}\frac{\partial f}{\partial\bar{\delta}\tau}+i\delta\bar{\delta}\check{\omega}\frac{\partial f}{\partial\delta\tau}-i\bar{\delta}\delta\check{\omega}\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$
The equation must be satisfied for any values of $\check{\kappa},\check{\lambda},\check{\omega}$, so that we obtain
$2D\rho\frac{\partial f}{\partial D\rho}+2D\sigma\frac{\partial f}{\partial D\sigma}+2D\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}+\delta\rho\frac{\partial f}{\partial\delta\rho}+\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\rho}+\delta\sigma\frac{\partial f}{\partial\delta\sigma}$
$+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+\rho\frac{\partial f}{\partial\rho}+\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}=0,$ (2.8v)
$\rho\frac{\partial f}{\partial D\rho}+\sigma\frac{\partial f}{\partial D\sigma}+\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}=0,$ (2.8w)
$-2D\sigma\frac{\partial f}{\partial D\sigma}+2D\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}+\delta\rho\frac{\partial f}{\partial\delta\rho}-\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\rho}-\delta\sigma\frac{\partial f}{\partial\delta\sigma}-3\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}$
$+3\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8x)
$-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8y)
$-\frac{\partial f}{\partial\tau}+2\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}-2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8z)
$\rho\frac{\partial f}{\partial\delta\rho}+\sigma\frac{\partial f}{\partial\delta\sigma}+\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8aa)
$\rho\frac{\partial f}{\partial\bar{\delta}\rho}+\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8ab)
$D\rho\frac{\partial f}{\partial\delta\rho}+D\sigma\frac{\partial f}{\partial\delta\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}-\sigma\frac{\partial f}{\partial\tau}+\rho\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8ac)
$D\rho\frac{\partial f}{\partial\bar{\delta}\rho}+D\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+\rho\frac{\partial f}{\partial\tau}-\bar{\sigma}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8ad)
$\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\bar{\tau}}=0,~\frac{\partial f}{\partial\delta\tau}=0,~\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}=0.$ (2.8ae)
Let us first assume $\rho\neq 0,|\sigma|\neq 0$. We have 9 partial differential equations $F^{i}f=0$ for $f$, depending on 14 independent variables $D\rho,D\sigma,D\bar{\sigma},\delta\rho,\bar{\delta}\rho,\delta\sigma,\bar{\delta}\sigma,\delta\bar{\sigma},\bar{\delta}\bar{\sigma},\rho,\sigma,\bar{\sigma},\tau,\bar{\tau}.$ The 36 integrability conditions $F^{i}(F^{k})f-F^{k}(F^{i})f=0$ are satisfied: They are identical to $F^{i}f=0$, as shown by a direct calculation. The system can be resolved with respect to the first $m=9$ derivatives of $f$, $\frac{\delta f}{D\rho},...\frac{\delta f}{\bar{\delta}\bar{\sigma}}$. One thus obtains its so-called canonical form. However, it should be noted that some denominators may become zero in this process, meaning that singular cases must be considered. All resolved derivatives contain the factor
$\rho|\sigma|I_{2}$ (2.8af)
in their denominators. Here $I_{2}$ is the imaginary part of the complex quantity
$I=I_{1}+iI_{2}=\frac{i}{|\sigma|}\Bigl(\frac{D\rho}{\rho}-\frac{D\sigma}{\sigma}\Bigr)+2\frac{\nu}{|\sigma|}.$ (2.8ag)
II is itself an invariant of the system, written here with a non-vanishing rotation coefficient $\nu$. To simplify the calculations, we assumed $\nu=0$. In the invariants $J,L,M,N$ to be derived, $\nu$ only appears via $I$, thus no transformation to $\nu\neq 0$ is necessary.
Obviously, we have to consider the following cases separately:
$|\sigma|\neq 0,~\rho\neq 0,~I_{2}\neq 0,$ (2.8ah)
$|\sigma|\neq 0,~\rho\neq 0,~I_{2}=0,$ (2.8ai)
$|\sigma|\neq 0,~\rho=0,$ (2.8aj)
$|\sigma|=0,~\rho=0.$ (2.8ak)
We start the first case
$|\sigma|\neq 0,~\rho\neq 0,~I_{2}\neq 0$
The invariants $I_{1},I_{2},j$ may be used to find further invariants, following a method described by Kamke [18], p. 54. The independent variables $D\sigma,~D\bar{\sigma},~\rho$ can be replaced by the invariants $I_{1},I_{2},j$:
$D\sigma=(iI_{1}-I_{2})\sigma\sqrt{\sigma\bar{\sigma}}+D\rho\sqrt{\sigma}/(\sqrt{\bar{\sigma}}j),$ (2.8al)
$D\bar{\sigma}=-(iI_{1}+I_{2})\bar{\sigma}\sqrt{\sigma\bar{\sigma}}+D\rho\sqrt{\bar{\sigma}}/(\sqrt{\sigma}j),$ (2.8am)
$\rho=j\sqrt{\sigma\bar{\sigma}}.$ (2.8an)
The function $f$ therefore no longer depends on $D\sigma,D\bar{\sigma},\rho$. $I_{1},I_{2},j$ are to be regarded as parameters. We have to consider the new system
$\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}+\delta\rho\frac{\partial f}{\partial\delta\rho}+\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\rho}+\delta\sigma\frac{\partial f}{\partial\delta\sigma}$
$+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8ao)
$\frac{\partial f}{\partial D\rho}=0,$ (2.8ap)
$\delta\rho\frac{\partial f}{\partial\delta\rho}-\bar{\delta}\rho\frac{\partial f}{\partial\bar{\delta}\rho}-\delta\sigma\frac{\partial f}{\partial\delta\sigma}-3\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}$
$+3\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8aq)
$-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8ar)
$-\frac{\partial f}{\partial\tau}+2\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}-2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8as)
$j|\sigma|\frac{\partial f}{\partial\delta\rho}+\sigma\frac{\partial f}{\partial\delta\sigma}+\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8at)
$j|\sigma|\frac{\partial f}{\partial\bar{\delta}\rho}+\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8au)
$j^{2}|\sigma|^{2}\frac{\partial f}{\partial\bar{\tau}}-j|\sigma|^{2}e^{is}\frac{\partial f}{\partial\tau}+j|\sigma|D\rho\frac{\partial f}{\partial\delta\rho}$
$+\sigma(ij|\sigma|^{2}I_{1}-j|\sigma|^{2}I_{2}+D\rho)\frac{\partial f}{\partial\delta\sigma}$
$+\bar{\sigma}(-ij|\sigma|^{2}I_{1}-j|\sigma|^{2}I_{2}+D\rho)\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8av)
$j^{2}|\sigma|^{2}\frac{\partial f}{\partial\tau}-j|\sigma|^{2}e^{-is}\frac{\partial f}{\partial\bar{\tau}}+j|\sigma|D\rho\frac{\partial f}{\partial\bar{\delta}\rho}$
$+\sigma(ij|\sigma|^{2}I_{1}-j|\sigma|^{2}I_{2}+D\rho)\frac{\partial f}{\partial\bar{\delta}\sigma}$
$+\bar{\sigma}(-ij|\sigma|^{2}I_{1}-j|\sigma|^{2}I_{2}+D\rho)\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0$ (2.8aw)
for the new function $f=f(D\rho,\delta\rho,\bar{\delta}\rho,\delta\sigma,\bar{\delta}\sigma,\delta\bar{\sigma},\bar{\delta}\bar{\sigma},\sigma,\bar{\sigma},\tau,\bar{\tau}).$ The second equation shows that $f$ is also independent of $D\rho$. The system then consists of 8 equations for 10 independent variables $\delta\rho...\bar{\tau}$. An integral basis should therefore consist of 10-8=2 functions. The complex quantity
$J=J_{1}+iJ_{2}=e^{is/2}\Bigl[(\vartheta\bar{\vartheta}-4)\Bigl(\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}\Bigr)+(\zeta\bar{\vartheta}-4)\Bigl(\frac{\delta\bar{\sigma}}{\bar{\sigma}}-\frac{\delta\rho}{\rho}\Bigr)$
$+2\bar{\vartheta}(\zeta-\vartheta)\bar{\tau}\Bigr]+2e^{-is/2}(\zeta-\vartheta)\Bigl(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\rho}{\rho}+2\tau\Bigr)$ (2.8ax)
with
$\zeta=2j-iI_{1}+I_{2},~\vartheta=2j-iI_{1}-I_{2},~e^{is}=\sigma/|\sigma|$ (2.8ay)
is a solution and represents an integral basis.- In the case
$|\sigma|\neq 0,~\rho\neq 0,~I_{2}=0$,
the invariant $J$ simplifies to
$J=(I_{1}^{2}-4(1-j^{2}))M,$ (2.8az)
where
$M=e^{is/2}(\frac{\delta\sigma}{\sigma}+\frac{\delta\bar{\sigma}}{\bar{\sigma}}-2\frac{\delta\rho}{\rho})$ (2.8ba)
itself is a new complex invariant in the case $I_{2}=0$. As abbreviation we frequently use
$I_{3}=I_{1}^{2}-4(1-j^{2}).$ (2.8bb)
If also $I_{3}=0$ applies, there is a further invariant
$L=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}-2\bar{\tau})+e^{-is/2}(j+i\sqrt{1-j^{2}})(\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}-\frac{\bar{\delta}\rho}{\rho}-2\tau).$ (2.8bc)
The next special case to consider is non-vanishing shear with vanishing divergence:
$|\sigma|\neq 0,~\rho=0$
Here the following system must be solved;
$2D\sigma\frac{\partial f}{\partial D\sigma}+2D\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}+\delta\sigma\frac{\partial f}{\partial\delta\sigma}$
$+\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}+\sigma\frac{\partial f}{\partial\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}=0,$ (2.8bd)
$\sigma\frac{\partial f}{\partial D\sigma}+\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}=0,$ (2.8be)
$-2D\sigma\frac{\partial f}{\partial D\sigma}+2D\bar{\sigma}\frac{\partial f}{\partial D\bar{\sigma}}-\delta\sigma\frac{\partial f}{\partial\delta\sigma}-3\bar{\delta}\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}$
$+3\delta\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}+\bar{\delta}\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-2\sigma\frac{\partial f}{\partial\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\bar{\sigma}}-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8bf)
$-\frac{\partial f}{\partial\bar{\tau}}-2\sigma\frac{\partial f}{\partial\delta\sigma}+2\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8bg)
$-\frac{\partial f}{\partial\tau}+2\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}-2\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8bh)
$\sigma\frac{\partial f}{\partial\delta\sigma}+\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}=0,$ (2.8bi)
$\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}=0,$ (2.8bj)
$D\sigma\frac{\partial f}{\partial\delta\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\delta\bar{\sigma}}-\sigma\frac{\partial f}{\partial\tau}=0,$ (2.8bk)
$D\sigma\frac{\partial f}{\partial\bar{\delta}\sigma}+D\bar{\sigma}\frac{\partial f}{\partial\bar{\delta}\bar{\sigma}}-\bar{\sigma}\frac{\partial f}{\partial\bar{\tau}}=0,$ (2.8bl)
$\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\bar{\tau}}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8bm)
This time, there are 9 differential equations for 10 independent variables $D\sigma,D\bar{\sigma},\delta\sigma,\bar{\delta}\sigma,\delta\bar{\sigma},\bar{\delta}\bar{\sigma},\sigma,\bar{\sigma},\tau,\bar{\tau}$. The integrability conditions are also satisfied, as can be easily verified. When attempting to convert the equations into canonical form, again terms appear in some denominators that become zero in the special case $I_{1}=0$, equivalent to $Ds=0$. We ignore this case initially. The last equation in the canonical version can be written:
$\frac{\delta f}{\delta\bar{\delta}\bar{\sigma}}(4-I_{1}^{2})\frac{|\sigma|^{3}e^{-is}}{Ds}=0.$ (2.8bn)
In the case of $I_{1}^{2}=4$, this equation drops out, and the integration of the remaining equations yields the invariant
$N=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\bar{\sigma}}{\bar{\sigma}}-4\bar{\tau})/(1+i)-ie^{-is/2}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}+4\tau)/(1+i).$ (2.8bo)
The apparently complex function $N$ is actually real and represents one single invariant. If $I_{1}^{2}\neq 4$, the only invariant for this case is - apart from $j=0$ - $I_{1}$. This case also includes the special case $I_{1}=0$ above.
In the case $|\sigma|=0,~\rho\neq 0$ we are looking for invariants with dependencies $f=f(D\rho,\delta\rho,\bar{\delta}\rho,\delta\tau,\bar{\delta}\tau,\delta\bar{\tau},\bar{\delta}\bar{\tau},\rho,\tau,\bar{\tau})$. Using the same methods as before, it can be shown that no invariants exist.
The last case is $|\sigma|=0,~\rho=0$. The invariants depend here on the independent variables $\delta\tau,\bar{\delta}\tau,\delta\bar{\tau},\bar{\delta}\bar{\tau},\tau,\bar{\tau}$. One obtains the differential equations
$2\tau\frac{\partial f}{\partial\delta\tau}+\frac{\partial f}{\partial\bar{\tau}}=0,~\bar{\tau}(\frac{\partial f}{\partial\delta\tau}+\frac{\partial f}{\partial\bar{\delta}\bar{\tau}})+\frac{\partial f}{\partial\tau}=0,~-\tau\frac{\partial f}{\partial\tau}+\bar{\tau}\frac{\partial f}{\partial\bar{\tau}}=0,$~ (2.8bp)
$\frac{\partial f}{\partial\delta\bar{\tau}}=\frac{\partial f}{\partial\bar{\delta}\tau}=0,~\frac{\partial f}{\partial\delta\tau}=\frac{\partial f}{\partial\bar{\delta}\bar{\tau}}.$ (2.8bq)
A solution $f$ is the Gaussian curvature
$K=-2\tau\bar{\tau}+\bar{\delta}\bar{\tau}+\delta\tau.$
or any function of $K$.
Appendix $A$ provides a brief overview of the results of this section. One important point should be mentioned: There exist no unique set of invariants valid for all null hypersurfaces. Different classes of hypersurfaces admit different sets of differential invariants.
4 The intrinsic Killing equation for $\mathcal{N}_{3}$.
After these preparations we discuss Killing isometries of null hypersurfaces. A $\mathcal{N}_{3}$ admits an inner isometry group if the Lie derivative of the inner metric $\gamma_{ik}$ vanishes with respect to a Killing field $\xi^{i}(x^{i})$. The Lie derivative is already defined on a differentiable manifold and therefore independent of any affinity given additionally. It can be applied to degenerate metrics such as that of a null hypersurface, where an affinity is not uniquely determined:
$\mathrel{{\mathop{L}\limits_{\xi^{i}}}}\gamma_{ik}\equiv\frac{\partial\gamma_{ik}}{\partial x^{l}}\xi^{l}+\gamma_{kl}\frac{\partial\xi^{l}}{\partial x^{i}}+\gamma_{il}\frac{\partial\xi^{l}}{\partial x^{k}}.$ (2.8a)
But with the affinities (2.8i) the Lie derivative may also be written
$\mathrel{{\mathop{L}\limits_{\xi^{i}}}}\gamma_{ik}=\xi^{l}\nabla_{l}\gamma_{ik}+\gamma_{kl}\nabla_{i}\xi^{l}+\gamma_{il}\nabla_{k}\xi^{l},$ (2.8b)
where $\nabla_{i}$ is the covariant derivative associated with the affinity. With $\xi_{k}=\gamma_{kl}\xi^{l}$ and $\nabla_{i}\gamma_{kl}=\gamma_{k}h_{li}+\gamma_{l}h_{ki}$ the expression is transformed into a third form
$\mathrel{{\mathop{L}\limits_{\xi^{i}}}}\gamma_{ik}=\nabla_{k}\xi_{i}+\nabla_{i}\xi_{k}-2\xi^{l}\gamma_{l}h_{ik}.$ (2.8c)
$h_{ik}$ is the Lie derivative of $\gamma_{ik}$ with regard to $\epsilon^{l}$:
$h_{ik}=\mathrel{{\mathop{L}\limits_{\epsilon^{i}}}}\gamma_{ik}=\frac{1}{2}(\partial_{k}\gamma_{il}+\partial_{i}\gamma_{kl}-\partial_{l}\gamma_{ik})\epsilon^{l},$ (2.8d)
where $\partial_{k}$ is the ordinary derivative. Thus the intrinsic Killing equation (KE) of a null hypersurface can also be written in the form
$\nabla_{k}\xi_{l}+\nabla_{i}\xi_{k}=2h_{ik}\xi^{l}\gamma_{l},$ (2.8e)
which differs from the usual covariant KE in non-degenerated metrics by an addtional term on the rhs.
We now ask which groups of motions on null hypersurfaces exist and which metrics are compatible with the existence of motions. The discussion is limited to local areas, where a Killing vector (KV) $\xi^{i}$ points either everywhere or nowhere in the direction of the generators. Likewise, the assumptions made below about the disappearance or non-disappearance of invariants or covariants shall always apply to a finite subdomain.
In this section, the triad formalism is tentatively used to answer the questions. In general, we can represent $\xi^{i}$ as
$\xi^{i}=\xi t^{i}+\bar{\xi}\bar{t}^{i}+\eta\epsilon^{i}.$ (2.8f)
The KE can then be written
$\bar{\delta}\xi+\xi\tau-\eta\sigma$ $=$ $0,$ (2.8g)
$\delta\xi+\bar{\delta}\bar{\xi}-\xi\bar{\tau}-\bar{\xi}\tau-2\rho\eta$ $=$ $0,$ (2.8h)
$D\xi+(\rho+i\nu)\xi+\sigma\bar{\xi}$ $=$ $0.$ (2.8i)
Obviously there exist two types of solutions of the KE: $\xi^{i}$ can point in the generator (null) direction (hence $\xi=0$) or in space-like directions ($\xi\bar{\xi}\neq 0$). In each case we have to study the integrability conditions of the KE. Their discussion requires some case distinctions. We assume that the conditions for each case are always valid for a finite domain and start with the general case $\rho\neq 0,~\sigma\neq 0,~I_{2}\neq 0$. (2.8h) permits to eliminate $\eta$:
$\eta=\frac{1}{2\rho}(\delta\xi+\bar{\delta}\bar{\xi}-\xi\bar{\tau}-\bar{\xi}\tau).$ (2.8j)
(2.8g) then takes the form
$\bar{\delta}\xi+\xi\tau-\frac{\sigma}{2\rho}(\delta\xi+\bar{\delta}\bar{\xi}-\xi\bar{\tau}-\bar{\xi}\tau)=0.$ (2.8k)
[…]
$\delta\xi(1-\frac{iI}{2j})-\bar{\delta}\bar{\xi}(1+\frac{iI}{2j})=A\xi+B\bar{\xi},$ (2.8l)
$A=\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}-\bar{\tau}(1+\frac{iI}{2j}),~B=\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\rho}{\rho}+\tau(1-\frac{iI}{2j}).$ (2.8m)
Equation (2.8l) is a system of algebraic equations for $\delta\xi,\bar{\delta}\bar{\xi}$. If the determinant
$\frac{i}{j}(\bar{I}-I)=\frac{2I_{2}}{j}\neq 0,$ (2.8n)
i.e. if the invariant $I_{2}$ is different from zero as supposed, the solution is
$\delta\xi=R\xi+S\bar{\xi}$ (2.8o)
with
$R$ $=$ $\frac{j}{2I_{2}}(\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho})(1+\frac{i\bar{I}}{2j})+\frac{j}{2I_{2}}(\frac{\delta\bar{\sigma}}{\bar{\sigma}}-\frac{\delta\rho}{\rho})(1+\frac{iI}{2j}),$ (2.8p)
$S$ $=$ $\frac{j}{2I_{2}}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\rho}{\rho})(1+\frac{i\bar{I}}{2j})+\frac{j}{2I_{2}}(\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}-\frac{\bar{\delta}\rho}{\rho})(1+\frac{iI}{2j})+\tau.$ (2.8q)
Altogether, instead of (2.8g)-(2.8i), we get the system (2.8o) and
$\bar{\delta}\xi$ $=$ $\frac{\sigma\xi}{2\rho}(R+\bar{S}-\bar{\tau})-\xi\tau+\frac{\sigma\bar{\xi}}{2\rho}(S+\bar{R}-\tau),$ (2.8r)
$D\xi$ $=$ $-(\rho+i\nu)\xi-\sigma\bar{\xi}$ (2.8s)
We have to add the equation giving the Killing vector component in the null direction, $\eta$:
$\eta=\frac{1}{2\rho}\xi(R+\bar{S}-\bar{\tau})+\frac{1}{2\rho}\bar{\xi}(\bar{R}+S-\tau).$ (2.8t)
The equations (2.8o),(2.8r),(2.8s) are a system of linear differential equations for the complex function $\xi$. For such systems a well-known theorem applies [14], shortly ”Eisenharts theorem”: The system is completely integrable, if the integrability conditions are satisfied identically. One obtains a first set of integrability conditions from the commutator relations (2.8u)-(2.8w) applied to $\xi$:
$D\delta\xi-\delta D\xi=(\rho+i\upsilon)\delta\xi+\bar{\sigma}\bar{\delta}\xi,$ (2.8u)
$D\bar{\delta}\xi-\bar{\delta}D\xi=(\rho-i\upsilon)\bar{\delta}\xi+\sigma\delta\xi,$ (2.8v)
$\delta\bar{\delta}\xi-\bar{\delta}\delta\xi=\bar{\tau}\bar{\delta}\xi-\tau\delta\xi.$ (2.8w)
If we substitute the derivatives of $\xi$, (2.8u) leads to
$\xi L_{1}+\bar{\xi}M_{1}=0$ (2.8x)
with
$L_{1}$ $=$ $-\delta(\rho+i\nu)-DR+\bar{\sigma}S+(\rho+i\upsilon)R-\bar{\sigma}\tau,$ (2.8y)
$M_{1}$ $=$ $-\delta\sigma-DS+\sigma R+(\rho-i\upsilon)S+\sigma\bar{\tau}.$ (2.8z)
(2.8x) is a homogeneous system of algebraic equations for $\xi,\bar{\xi}$. It has the algebraic solvability condition
$L_{1}\bar{L}_{1}-M_{1}\bar{M}_{1}=0.$ (2.8aa)
This condition is necessary for the existence of a Killing symmetry on $\mathcal{N}_{3}$ in the general case $\rho\neq 0,~\sigma\neq 0,I_{2}\neq 0$. Suppose it holds, we have to distinguish three cases. First, let
$M_{1}\bar{M}_{1}+L_{1}\bar{L}_{1}\neq 0,~M_{1}+\bar{L}_{1}\neq 0.$
The solution in this case is
$\xi=iw(\bar{L}_{1}+M_{1})$ (2.8ab)
with real $w$. In the case
$M_{1}\bar{M}_{1}+L_{1}\bar{L}_{1}\neq 0,~M_{1}+\bar{L}_{1}=0$
the solution can be written
$\xi=wM_{1}$ (2.8ac)
again with real $w$. In the third case $L_{1}=M_{1}=0$ the inegrability conditions are satisfied without restriction for $\xi$.
The two other equations (2.8v), (2.8w) lead to similar homogeneous algebraic equations for $\xi$, with still more complicated invariant coefficients $L_{2},M_{2},M_{3},M_{3}$, and to additional necessary conditions for the existence of a spacelike Killing symmetry:
$L_{2}\bar{L}_{2}-M_{2}\bar{M}_{2}=0,~L_{3}\bar{L}_{3}-M_{3}\bar{M}_{3}=0.$ (2.8ad)
Therefore, in the general case $\rho\neq 0,~\sigma\neq 0,~I_{2}\neq 0$ the integrability conditions split into a large number of cases. They are difficult to handle without an elaborated theory of higher-order null hypersurface invariants, which does not yet exist. We will therefore use the coordinate formalism for further calculations. Special cases like null Killing vectors can still be treated with the triad formalism.
5 Groups of motion with a null Killing vector
5.1 Basic equations
Consider the case that a Killing vector $\mathrel{{\mathop{\xi}\limits}}^{i}$ points to the null direction $\epsilon^{i}$, hence $\xi=0$ in the representation (2.8b). From the Killing equations (2.8g)-(2.8i) we conclude that the divergence $\rho$ and shear $\sigma$ vanish. Equivalently, the Lie derivative of the metric with regard to $\epsilon^{i}$ vanishes, $h_{ik}=0$.
This type of null hypersurface is called Killing horizon (or simply horizon). Physically there are many different types of horizons, depending on their embedding in ambient spacetime. But from the point of view of inner geometry they all reduce to a null hypersurface with vanishing shear and divergence.
The Lie group of a horizon represents an infinite Lie group $G_{\infty}$ rather than a one-dimensional group, since with $X=\xi^{k}\partial_{k}$ every $X^{\prime}=f(x^{i})X$ is also a group generator, $f(x^{i})\neq 0$ is an otherwise arbitrary function of position. The metric of a general Killing horizon is in adapted coordinates:
$\mathbb{G_{\infty}}:~~~E=E(x_{2},x_{3}),~F=F(x_{2},x_{3}),~G=G(x_{2},x_{3}).$ (2.8a)
This is the metric of a general two-dimensional surface. Its Gaussian curvature $K$ (see (2.8ah)) is the only second-order differential invariant of horizons.
Horizons may admit additionally motions with spacelike trajectories. The KEs for horizons simplify to
$\bar{\delta}\xi+\xi\tau$ $=$ $0,$ (2.8b)
$\delta\xi+\bar{\delta}\bar{\xi}-\xi\bar{\tau}-\bar{\xi}\tau$ $=$ $0,$ (2.8c)
$D\xi+i\nu\xi$ $=$ $0.$ (2.8d)
Obviously, $\xi=0$ is the horizon solution. We look for additional solution with $|\xi\bar{\xi}|\neq 0$.
In order to apply Eisenhart’s theorem, the form of (2.8b)-(2.8d) is not yet suitable: Not all first-order derivatives can be expressed by the functions themselves. However, this is the case for the equivalent system:
$\delta h$ $=$ $\xi\delta\bar{\tau}+\bar{\tau}h+K\bar{\xi},$ (2.8e)
$\bar{\delta}h$ $=$ $\xi(\tau\bar{\tau}-\delta\tau),$ (2.8f)
$Dh$ $=$ $-i\xi\delta\upsilon,$ (2.8g)
$\delta\xi$ $=$ $h,$ (2.8h)
$\bar{\delta\xi}$ $=$ $-\xi\tau,$ (2.8i)
$D\xi$ $=$ $-i\upsilon\xi$ (2.8j)
with the constraint
$h+\bar{h}=\xi\bar{\tau}+\bar{\xi}\tau.$ (2.8k)
$K$ is the quantity defined by (2.8ah). If the integrability conditions for the extended system are satisfied identically, the general solution would involve 4 arbitrary constants, but related by one constraint. Hence a 3-parameter group with space-like generators can be expected at most. A first set of integrability conditions is obtained by applying the D-operator on (2.8e)-(2.8k) and substitution the first derivatives:
$\xi DK$ $=$ $0,$ (2.8l)
$\bar{\xi}DK$ $=$ $0,$ (2.8m)
$(\bar{\delta}\bar{\xi}+\delta\xi)K+\xi(\delta K-K\bar{\tau})+\bar{\xi}({\bar{\delta}}K-K\tau)$ $=$ $0,$ (2.8n)
$D\bar{h}-i\bar{\xi}\bar{\delta}\upsilon$ $=$ $0,$ (2.8o)
$-\delta\bar{\xi}\tau+\bar{\xi}(\bar{\delta}\bar{\tau}-2\tau\bar{\tau})+\delta\bar{h}$ $=$ $0,$ (2.8p)
$-\bar{\delta}\bar{\xi}\tau-\xi K-\bar{\xi}\bar{\delta}\tau+\bar{\delta}\bar{h}$ $=$ $0.$ (2.8q)
Since $\xi\bar{\xi}\neq 0$, the first two equations require $DK=0$, thus $K$ cannot change in the null direction. Taking the constraint (2.8k) into account, the third equation can be written
$\xi\delta K+\bar{\xi}\bar{\delta}K=0.$ (2.8r)
This is again an algebraic equation for $\xi$, but with already satisfied solvability conditions.
5.2 $\delta K\neq 0$.
We assume $\delta K\neq 0$ and treat the case $\delta K=0$ later. Equation (2.8r) is then solved by
$\xi=2iw\bar{\delta}K$ (2.8s)
with a real function $w$. Introducing this ansatz in the original system (2.8b)-(2.8d), one obtains
$\bar{\delta}KDw=0,$ (2.8t)
$\delta w\delta K+w(\delta\delta K+\bar{\tau}\delta K)=0,$ (2.8u)
$\delta w\bar{\delta}K-\delta K\bar{\delta}w=0.$ (2.8v)
The first equation requires $Dw=0$, i.e. $w$ remains unchanged when progressing in the null direction. The change of $w$ in spatial directions is determined by the second equation, while the third one implies $Re(\delta w)Im(\delta K)=0$ and either restricts $\delta w$ or $\delta K$. We study these integrability conditions by writing them in coordinate form. Since $h_{ik}=0$, it follows from (2.8d) that the $\gamma_{AB}$ depend only on $x_{2},x_{3}$. The Killing equations (2.8a) read in coordinate form:
$\xi^{B}_{,1}=0,$ (2.8w)
$\gamma_{AB,C}\xi^{C}+\gamma_{BC}\xi^{C}_{,A}+\gamma_{AC}\xi^{C}_{,B}=0.$ (2.8x)
With help of coordinate transformations (2.8an), $\xi^{B}=\delta^{B}_{3}$ can be reached. The metric then satisfies $\gamma_{AB,3}=0$. The transformations that are still allowed form a subgroup of (2.8am),(2.8an):
$x^{\prime}_{1}=f(x_{1},x_{2}),~x^{\prime}_{2}=g(x_{2}),~x^{\prime}_{3}=x_{3}+h(x_{2})$ (2.8y)
with $f_{,1}g_{,2}\neq 0$. Using (2.8y), we may obtain
$ds^{2}=dx_{2}^{2}+m(x_{2})dx_{3}^{2}$ (2.8z)
with a single positive function $m(x_{2})$. The spatial components $t^{2},t^{3}$ of the triad may be chosen as $t^{2}=\frac{i}{\sqrt{2}},~t^{3}=\frac{1}{\sqrt{2m}}$. From (2.8f) and $\xi^{k}=\delta^{k}_{3}$ then follows
$\xi=\sqrt{m/2}.$ (2.8aa)
Comparison with (2.8s) yields
$w=\frac{m}{2K_{,2}}.$ (2.8ab)
$K$ calculated for the metric (2.8z) can be written
$K=(-2mm_{,22}+m_{,2}^{2})/m^{2}$ (2.8ac)
and depends on $x_{2}$ only. Since $|\delta K|\neq 0$ was assumed, we have $K_{,2}\neq 0$.
Discussing the integrability conditions (2.8t)-(2.8v) becomes now trivial. (2.8t) is obviously satisfied. Since $Re~\delta w=0$, (2.8v) holds. A short calculation shows that also (2.8u) is satified. We have a $G_{1}$ as isometry group with a spacelike generator, together with a $G_{\infty}$:
$\mathbb{G_{\infty}x~G_{1}I}:~~~E=1,~F=0,~G=m(x_{2}).$ (2.8ad)
5.3 $\delta K=0$
In the case $\delta K=0$ the equation (2.8r) is trivially solved. Together with $DK=0$, the invariant $K$ is then a constant on the null surface. The remaining three integrability conditions are also satisfied, if the initial extended system is considered. Hence, as in the case $\delta K\neq 0$, also for horizons with $K=const$ all integrability conditions of the extended system are satisfied identically. From Eisenhart’s theorem one expects at most a $G_{3}$ group of spatial motions, together with the $G_{\infty}$.
All these facts are well-known from the differential geometry of two-dimensional surfaces (cf. [23]). With the coordinate transformations (2.8an) we may obtain normal (or canonical) forms for the metric $(a=const)$, we list also the corresponding KVs:
$K=0$
$\mathbb{G_{\infty}x~G_{3}VII_{0}}:E=a^{2},~F=0,~G=a^{2}$ (2.8ae)
$\mathrel{{\mathop{\xi}\limits_{0}}}^{i}=(f(x_{1}),0,0),~\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=(0,0,1),~\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(0,1,0),~\mathrel{{\mathop{\xi}\limits_{3}}}=(0,x_{3},-x_{2}).$ (2.8af)
$K=-1/a^{2}$
$\mathbb{G_{\infty}x~G_{3}VIII}:E=a^{2}/x_{2}^{2},~F=0,~G=a^{2}/x_{2}^{2}$ (2.8ag)
$\mathrel{{\mathop{\xi}\limits_{0}}}^{i}=(f(x_{1}),0,0),~\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=(0,0,1),~\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(0,x_{2},x_{3}),$ (2.8ah)
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(0,2x_{2}x_{3},x_{3}^{2}-x_{2}^{2}).$
$K=1/a^{2}$
$\mathbb{G_{\infty}x~G_{3}IX}:E=a^{2},~F=0,~G=a^{2}\cos{x_{2}}^{2}$ (2.8ai)
$\mathrel{{\mathop{\xi}\limits_{0}}}^{i}=(f(x_{1}),0,0),~\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=(0,0,1),~\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(0,\cos{x_{3}},\sin{x_{3}}\tan{x_{2}}),$ (2.8aj)
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(0,-\sin{x_{3}},\cos{x_{3}}\tan{x_{2}}).$
The $G_{3}$ part of the motion group is classified according to Bianchi, $G_{3}VII_{0}$ is the Bianchi type $G_{3}VII$ with $q=0$, cf. (7.1). A horizon metric does not always appear in the canonical form given above. Frequently, the metric
$E=a,~F=ce^{-x_{2}},~G=be^{-2x_{2}}$ (2.8ak)
($a,b,c=const$) arises as limiting case of space-like $G_{2}$, $G_{3}$ and $G_{4}$ symmetries. Actually, (2.8ak) is a horizon: Using (2.8y) with
$x^{\prime}_{2}=e^{x_{2}}\sqrt{ab-c^{2}}/b,~x^{\prime}_{3}=x_{3}+ce^{x_{2}}/b$ (2.8al)
we obtain the normal form of a $G_{\infty}\times G_{3}VIII$ metric with
$E^{\prime}=a^{\prime 2}/x_{2}^{\prime 2}=(ab-c^{2})/(bx_{2}^{\prime}).$ (2.8am)
5.4 Example: horizons of the Kerr metric
An example of horizons of the type $G_{\infty}\times G_{1}$ is provided by the Kerr metric, which in Finkelstein-Kruskal coordinates can be written
$ds^{2}=-(1-\frac{2mr}{\Sigma})dt^{2}+2dtdr-\frac{4amrsin^{2}\theta}{\Sigma}drd\phi-2asin^{2}\theta drd\phi$ (2.8an)
$+\Sigma d\theta^{2}+(\frac{(r^{2}+a^{2})^{2}-a^{2}\sin^{2}\theta\Delta}{\Sigma})\sin^{2}\theta d\phi$
with
$\Sigma=r^{2}+a^{2}\cos^{2}\theta,~\Delta=r^{2}-2mr+a^{2}.$ (2.8ao)
$\Delta$ has two zeros $r_{-}$ and $r_{+}$. The equation $r=r_{\pm}$ represents the two horizons. From (2.8an) one obtains the induced metric of the horizons:
$\gamma_{11}=\frac{a^{2}\sin^{2}\theta}{\Sigma_{\pm}},~\gamma_{12}=0,~\gamma_{13}=-\frac{2amr_{\pm}}{\Sigma_{\pm}}\sin^{2}{\theta},$ (2.8ap)
$\gamma_{22}=\Sigma_{\pm},~~~~~~~~\gamma_{23}=0,~\gamma_{33}=\frac{4m^{2}r_{\pm}^{2}}{\Sigma_{\pm}}\sin^{2}{\theta}.$
A direct calculation gives indeed $det|\gamma_{ik}|=0.$ A suitable triad in non-adapted coordinates can be found as
$\epsilon^{i}=(2am,0,2m-r_{\pm}),$ (2.8aq)
$t^{i}=(0,\frac{i}{\sqrt{2\Sigma_{\pm}}},\frac{\sqrt{\Sigma_{\pm}}}{2\sqrt{2}mr_{\pm}\sin{\theta}}),$ (2.8ar)
$\gamma_{i}=(\frac{1}{2am},0,0),$ (2.8as)
$t_{i}=(-\frac{a\sin{\theta}}{\sqrt{2\Sigma_{\pm}}},i\sqrt{\frac{\Sigma_{\pm}}{2}},\sqrt{\frac{2}{\Sigma_{\pm}}}mr_{\pm}\sin{\theta}).$ (2.8at)
The transformation
$x^{\prime}_{1}=x_{1},~x^{\prime}_{2}=x_{2},~x^{\prime}_{3}=x_{1}/(2am)-x_{3}/(2m-r_{\pm})$ (2.8au)
($x_{2}\equiv\theta,~x_{3}\equiv\phi$) leads to adapted coordinates with $\gamma^{\prime}_{11}=0,\gamma^{\prime}_{1A}=0$ and to
$\gamma^{\prime}_{22}=\Sigma_{\pm},$ (2.8av)
$\gamma^{\prime}_{23}=0,$~ (2.8aw)
$\gamma^{\prime}_{33}=\frac{4m^{2}a^{4}\sin^{2}{\theta}}{\Sigma_{\pm}}.$ (2.8ax)
In adapted coordinates the calculation of $K$ is easy, on obtains
$K=\frac{(a^{2}+r_{\pm}^{2})(r_{\pm}^{2}-3a^{2}\cos^{2}{\theta})}{(r_{\pm}^{2}+a^{2}\cos^{2}{\theta})^{3}},$ (2.8ay)
a formula first derived by Smarr [32].
6 Groups $G_{1},G_{2}$ with spacelike generators
We now study groups of motions $G_{r}$ with $r$ space-like generators $X=\mathrel{{\mathop{\xi}\limits_{k}}}^{i}\partial_{i},k=1...r.$ For the commutators we use Bianchi’s notation
[…]
The maximal number of multiple KS’s in a n-dimensional manifold is $n(n+1)/2$, therefore 6 for the $\mathcal{N}_{3}.$ A $G_{6}$ is not possible, unlike to a nondegenerate $V_{3}$. Fubinis theorem [16] also excludes a $G_{5}$. The reason for the reduced mobility for a $\mathcal{N}_{3}$ is the existence of the null direction. It causes an inherent asymmetry in null hypersurfaces.
In this section groups $G_{1},G_{2}$ are considered. We use adapted coordinates with the line element (2.8al). This shape of the metric is preserved under the transformations (2.8am),(2.8an). The Killing equation (2.8e) can be written in these coordinates
$\xi_{,1}^{B}=0,$ (2.8b)
$\xi^{1}\gamma_{AB,1}+\gamma_{AB,C}\xi^{C}+\gamma_{AC}\xi^{C}_{,B}+\gamma_{BC}\xi^{C}_{,A}=0.$ (2.8c)
Coordinates may further be chosen such that $\xi^{i}=\delta^{i}_{3}$, say. The metric is then restricted by $\gamma_{AB,3}=0$. The coordinate transformations preserving this condition as well as the chosen normal form for $\xi^{i}$ are given by (2.8y). If only one space-like Killing vector exists, we have a motion $G_{1}$ with the metric
$\mathbb{G_{1}:}~E=E(x_{1},x_{2}),~F=F(x_{1},x_{2}),~G=G(x_{1},x_{2}).$ (2.8d)
6.1 Abelian groups $G_{2}I$ with space-like generators
[…]
requires $\mathrel{{\mathop{\xi}\limits_{2}}}^{k}=\mathrel{{\mathop{\xi}\limits_{2}}}^{k}(x_{1},x_{2})$. The KE (2.8b),(2.8c) admit to write $\mathrel{{\mathop{\xi}\limits_{2}}}^{1}=\alpha(x_{1},x_{2}),~\mathrel{{\mathop{\xi}\limits_{2}}}^{2}=\beta(x_{2}),~\mathrel{{\mathop{\xi}\limits_{2}}}^{3}=\gamma(x_{2})$, and lead to
$\alpha E_{,1}+\beta E_{,2}+2E\beta_{,2}+2F\gamma_{,2}$ $=$ $0,$
$\alpha F_{,1}+\beta F_{,2}+F\beta_{,2}+G\gamma_{,2}$ $=$ $0,$ (2.8f)
$\alpha G_{,1}+\beta G_{,2}$ $=$ $0.$
The functions $\alpha,\beta,\gamma$ transform under (2.8y) as
$\alpha^{\prime}$ $=$ $\alpha f_{,1}+\beta f_{,2},$
$\beta^{\prime}$ $=$ $\beta g_{,2},$ (2.8g)
$\gamma^{\prime}$ $=$ $\gamma+\beta h_{,2}.$
Two cases must be distinguished. Suppose first $\beta\neq 0$ (the case $\beta=0$ is treated in the following subsection). Then, by means of (6.1), $\beta=1,\gamma=0,\alpha=0$ can be reached, that is, $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=\delta_{2}^{i}$. The residual transformations
$x^{\prime}_{1}=f(x_{1}),~x^{\prime}_{2}=x_{2}+const,~x^{\prime}_{3}=x_{3}+const$ (2.8h)
with $f_{,1}\neq 0$ leave $\alpha,\beta$ and $\gamma$ unchanged. Integrating the resulting differential equation (6.1) gives a first normal form for a metric admitting a $G_{2}$ with space-like trajectories:
$\mathbb{G_{2}I_{.1}}:~~~E=a(x_{1}),~F=c(x_{1}),~G=b(x_{1}).$ (2.8i)
The rotation coefficients of this metric are
$~\rho=-d_{,1}/(4d),$ (2.8j)
$~\sigma=(a_{,1}b^{2}-b_{,1}ab+2b_{,1}c^{2}-2c_{,1}bc)/(4bd)+i(c_{,1}b-b_{,1}c)/(2b\sqrt{d}),$ (2.8k)
$~\nu=(b_{,1}c-c_{,1}b)/(2b\sqrt{d}),~\tau=0.$ (2.8l)
with $d=ab-c^{2}$. The differential invariants can be written if $|\sigma|=\sqrt{p}/(4d)\neq 0$:
$j$ $=$ $-d_{,1}/\sqrt{p},$ (2.8m)
$I_{1}$ $=$ $-8(d/p)^{3/2}(a_{,11}[c_{,1}b-cb_{,1}]+c_{,11}[b_{,1}a-ba_{,1}]$ (2.8n)
$+b_{,11}[a_{,1}c-ac_{,1}]),$
$I_{2}$ $=$ $j_{,1}/\rho,$ (2.8o)
$J$ $=$ $0$ (2.8p)
with
$p=(a_{,1}b-ab_{,1})^{2}+4(a_{,1}c-ac_{,1})(b_{,1}c-bc_{,1}).$ (2.8q)
Two special cases must be considered. In the case $|\sigma|=\sqrt{p}/(4d)=0$ (or $p=0$) and $\rho\neq 0$ (or $d_{,1}\neq 0$), the same metric applies (now denoted $\mathbb{G_{2}I_{.2}}$), but no invariants exist. If additionally $d=const$, this results in a constant metric with an additional space-like KV, a case treated later. Residual coordinate transformations are (2.8h).
6.2 Abelian groups $G_{2}I$ with spacelike generators coplanar with the null direction
If $\beta=0$, the situation is slightly more complicated. We suppose that $\alpha\neq 0$, otherwise $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ would be equivalent to $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}$ and no $G_{2}$ exists. Then, with $\alpha^{\prime}=\alpha f_{,1}$, we reach $\alpha^{\prime}=1$, and the simplified equation (6.1) can be integrated with the result:
$E$ $=$ $E_{0}(x_{2})-2\gamma_{,2}x_{1}F_{0}(x_{2})+\gamma_{,2}^{2}x_{1}^{2}G_{0}(x_{2}),$
$F$ $=$ $F_{0}(x_{2})-\gamma_{,2}x_{1}G_{0}(x_{2}),$ (2.8r)
$G$ $=$ $G_{0}(x_{2}).$
Since $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ must be spacelike, we have $\gamma(x_{2})\neq 0$. The still available coordinate freedom is
$x^{\prime}_{1}$ $=$ $x_{1}+f(x_{2}),$
$x^{\prime}_{2}$ $=$ $g(x_{2}),$ (2.8s)
$x^{\prime}_{3}$ $=$ $x_{3}+h(x_{2})$
with three free functions $f,g,h$ of $x_{2}$. A little calculation shows that under (6.2) $E_{0},F_{0},G_{0}$ change as
$g_{,2}^{2}E^{\prime}_{0}$ $=$ $E_{0}+2F_{0}(f\gamma_{,2}-h_{,2})+G_{0}(f\gamma_{,2}-h_{,2})^{2},$
$g_{,2}F^{\prime}_{0}$ $=$ $F_{0}+G_{0}(f\gamma_{,2}-h_{,2}),$ (2.8t)
$G^{\prime}_{0}$ $=$ $G_{0}$
Also, transforming the Killing vector $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(1,0,\gamma(x_{2}))$ leads to $\gamma^{\prime}(x_{2}^{\prime})=\gamma(x_{2})$. We can use these transformations to obtain a certain normal form of the metric. e.g., $E_{0}^{\prime}=1,~F^{\prime}_{0}=0$. Assuming $\gamma_{,2}\neq 0$, we choose
$f=\frac{1}{\gamma_{,2}}(h_{,2}-F_{0}/G_{0}),~g_{,2}^{2}=(E_{0}G_{0}-F_{0}^{2})/G_{0}$ (2.8u)
to obtain
$\mathbb{G_{2}I_{.3}}:$ (2.8v)
$E$ $=$ $1+x_{1}^{2}\gamma_{,2}^{2}m(x_{2}),$
$F$ $=$ $x_{1}\gamma_{,2}m(x_{2}),$ (2.8w)
$G$ $=$ $m(x_{2}).$
The still available coordinate freedom
$x^{\prime}_{1}=x_{1}-h_{,2}/\gamma_{,2},~x^{\prime}_{2}=x_{2}+const,~x^{\prime}_{3}=x_{3}+h(x_{2})$ (2.8x)
does not change this metric. The rotation coefficients are
$\rho=0,~|\sigma|=\frac{\sqrt{m}\gamma_{,2}}{2},$ (2.8y)
$\sigma=i\frac{\sqrt{m}\gamma_{,2}}{2},~\nu=-\frac{\sqrt{m}\gamma_{,2}}{2},~\tau=i\frac{m_{,2}}{2\sqrt{2}m}$ (2.8z)
For the invariants we obtain
$j=0,~I_{1}=-2\gamma_{,2}/|\gamma_{,2}|=-2sgn(\gamma_{2}),$ (2.8aa)
$N=(cos(s)-sin(s))m_{,2}/(\sqrt{2}m).$ (2.8ab)
The surfaces $x_{1}=const$ have the Gaussian curvature
$K=\frac{-2mm_{,22}+m_{,2}^{2}}{4m^{2}}.$ (2.8ac)
It is noteworthy that the metric (2.8v) allows a continous number of spacelike Killing fields commuting with the first two:
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(C_{,2}/\gamma_{,2},0,C),$ (2.8ad)
where $C$ is an arbitrary function of $x_{2}$. $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ is included for $C=\gamma$.
We treat now the case $\gamma=const$. Here the transformation functions $h,g$ satisfying
$h_{,2}=F_{0}/G_{0},~g_{,2}=\sqrt{(E_{0}G_{0}-F_{0}^{2})/G_{0}}$ (2.8ae)
result in
$E=1,~F=0,~G=m(x_{2}).$ (2.8af)
The residual transformations
$x^{\prime}_{1}=x_{1}+f(x_{2}),~x^{\prime}_{2}=x_{2},~x^{\prime}_{3}=x_{3}+const$ (2.8ag)
leave (2.8af) invariant. (2.8af) is the metric of the horizon (2.8ad).
The two types of Abelian groups $G_{2}I$ with space-like generators discussed here and in the previous subsection can be characterized by the quotient
$Q=\frac{\mathrel{{\mathop{\xi}\limits_{1}}}^{i}\mathrel{{\mathop{\xi}\limits_{2}}}^{k}\gamma_{ik}}{\sqrt{\mathrel{{\mathop{\xi}\limits_{1}}}^{i}\mathrel{{\mathop{\xi}\limits_{1}}}^{k}\gamma_{ik}}\sqrt{\mathrel{{\mathop{\xi}\limits_{2}}}^{i}\mathrel{{\mathop{\xi}\limits_{2}}}^{k}}\gamma_{ik}}.$ (2.8ah)
In the first case ($\beta\neq 0$) we have in adapted coordinates
$Q=\gamma_{23}/\sqrt{|\gamma_{AB}|+\gamma^{2}_{23}}<1,$ (2.8ai)
in the second case ($\beta=0$)
$Q=1.$ (2.8aj)
$Q$ is the cosinus of the angle between the space-like directions $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}$ and $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$, and $Q=1$ tells us that this angle is zero. Nevertheless, the two KVs represent different directions, they form a so-called null angle with each other [^2] The concept of null angles was introduced by R. Penrose in his Les Houches lectures 1973 [28].. The geometrical interpretation is that the two space-like directions span a light-like surface element, i.e. that the null direction $\epsilon^{i}$ can be represented as a linear combination of the KVs. In this case the vectors are called coplanar with the null direction. The criterion for two vectors $a^{i},b^{i}$ to be coplanar with the null direction $\epsilon^{i}$ is
$\epsilon_{ikl}a^{i}b^{k}\epsilon^{l}=0,$ (2.8ak)
where $\epsilon_{ikl}$ is the totally antisymmetric Levi-Civita symbol with $\epsilon_{123}=1$.
6.3 Non-Abelian groups $G_{2}II$ with spacelike generators
Non-Abelian groups of order 2 are denoted by $G_{2}II$ and satify the commutator relation
$X_{1},X_{2}]=X_{1}.$ (2.8al)
If the first KV is transformed into $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta_{3}^{i}$, this relation gives
$\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(\alpha(x_{1},x_{2}),\beta(x_{2}),x_{3}+\chi(x_{2}))$ (2.8am)
with a new function $\chi(x_{2})$. The transformations (2.8y) change $\alpha,\beta,\chi$:
$\alpha^{\prime}=f_{,1}\alpha+f_{,2}\beta,~\beta^{\prime}=g_{,2}\beta,~\chi^{\prime}=\chi-h+h_{,2}\beta.$ (2.8an)
Again two cases must be distinguished. Assuming first $\beta\neq 0$, we may reach $\alpha^{\prime}=0,\beta^{\prime}=1,\chi^{\prime}=0$. Integrating the KE for $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ yields the metric
$\mathbb{G_{2}II_{.1}:}~E=a(x_{1}),~F=c(x_{1})e^{-x_{2}},~G=b(x_{1})e^{-2x_{2}}.$ (2.8ao)
Still allowed are the coordinate transformations
$x_{1}^{\prime}=f(x_{1}),~x_{2}^{\prime}=x_{2}+g_{0},~x_{3}^{\prime}=x_{3}+h_{0}e^{x_{2}},$ (2.8ap)
where $f_{,1}\neq 0$ and an inverse function $x_{1}=\varphi(x_{1}^{\prime})$, $g_{0}$ and $h_{0}$ are constants. By means of (2.8ap) the metric (2.8ao) is transformed into
$E^{\prime}=a^{\prime}(x_{1}^{\prime})-2c^{\prime}(x_{1}^{\prime})h_{0}+b^{\prime}(x_{1}^{\prime})h_{0}^{2},~F^{\prime}=e^{g_{0}-x_{2}^{\prime}}(c^{\prime}(x_{1}^{\prime})-h_{0}b^{\prime}(x_{1}^{\prime})),~G^{\prime}=b^{\prime}(x_{1}^{\prime})e^{2(g_{0}-x_{2}^{\prime})}$ (2.8aq)
where $a^{\prime}(x_{1}^{\prime})\equiv a(\varphi(x_{1}^{\prime}))$ etc.
The rotation coefficients for the metric (2.8ao) are
$\rho=-d_{,1}/(4d),$ (2.8ar)
$\sigma=(a_{,1}b^{2}-b_{,1}ab+2b_{,1}c^{2}-2c_{,1}bc)/(4bd)+i(bc_{,1}-cb_{,1})/(2b\sqrt{d}),$ (2.8as)
$\nu=(b_{,1}c-c_{,1}b)/(2b\sqrt{d}),$ (2.8at)
$\tau=-i\sqrt{b/(2d)},~d\equiv ab-c^{2}.$ (2.8au)
The invariants can be written, if $d_{,1}\neq 0$:
$j=-d_{,1}/\sqrt{p},$ (2.8av)
$I_{1}=8(d/p)^{3/2}(a_{,11}(b_{,1}c-c_{,1}b)+b_{,11}(c_{,1}a-a_{,1}c)+c_{,11}(a_{,1}b-b_{,1}a)),$ (2.8aw)
$I_{2}=2\frac{d}{\sqrt{p}}(2\frac{d_{,11}}{d_{,1}}-\frac{p_{,1}}{p}),$ (2.8ax)
The function $p(x_{1})$ is again given by (2.8q). The invariants also include $J$, which, unlike $I$, is nonlinear in the second derivatives of the metric. $J$ describes the change of the geometry in the transverse directions and has a relatively complicated structure:
$J=-128e^{is/2}\frac{UV}{\sqrt{2b}p^{3}d_{,1}^{2}}$ (2.8ay)
with $e^{is/2}=\sqrt{\sigma/|\sigma|}$. $U=U_{1}+iU_{2}$ is complex with
$U_{1}=2bd^{2}d_{,1}(a_{,11}[bc_{1}-cb_{,1}]+b_{,11}[ca_{,1}-ac_{,1}]+c_{,11}[ab_{,1}-ba_{,1}])+dd_{1}p(ab_{,1}-ba_{,1}),$
$U_{2}=2bd^{5/2}(la_{,11}+mb_{,11}+2nc_{,11})$
$~~~~~~+pd^{1/2}(-b^{3}a_{,1}^{2}+b(c^{2}-3ab)a_{,1}b_{,1}+4b^{2}ca_{,1}c_{,1}$
$~~~~~~-ac^{2}b_{,1}^{2}+2c(c^{2}+ab)b_{,1}c_{,1}+2b(ab-3c^{2})c_{,1}^{2}).$
We have used the abbreviations:
$l=ba_{,1}b_{,1}-ab_{,1}^{2}+2cb_{,1}c_{,1}-2bc_{,1}^{2}$
$m=aa_{,1}b_{,1}-ba_{,1}^{2}+2ca_{,1}c_{,1}-2ac_{,1}^{2}$
$n=-2ca_{,1}b_{,1}+ab_{,1}c_{,1}+ba_{,1}c_{,1}.$
Finally, $V$ is a real function:
$V=d(la_{,11}+mb_{,,1}+nc_{,11})-p(a_{,1}b_{,1}-c_{,1}^{2})$
Again, two special cases must be considered.
In the case $d_{,1}=0$, the metric (2.8ao) admits - in addition to $j=0$ - only the invariant $I_{1}$:
$I_{1}=-sgn(c)\frac{2B}{\sqrt{d}A^{3/2}}$ (2.8az)
with
$A=(a_{,1}b+b_{,1}a)^{2}-4a_{,1}b_{,1}c^{2}$
$B=4dc^{2}(a_{,11}b_{,1}-b_{,,11}a_{,1})+ba_{,1}^{2}(a_{,1}b^{2}+b_{,1}ab-4b_{,1}c^{2})$
$-ab_{,1}^{2}(b_{,1}a^{2}+a_{,1}ab-4a_{,1}c^{2}).$
The surfaces $x_{1}=const$ have the constant negative curvature K=b/dK=-b/d.
The case $|\sigma|=0$ implies $p=0$ and leads to
$a(x_{1})=k_{1}b(x_{1}),~c(x_{1})=k_{2}b(x_{1})$ (2.8ba)
with the constants $k_{1},k_{2},~k_{1}>k_{2}^{2}$. We first assume $b_{,1}\neq 0.$ The metric is
$\mathbb{G_{2}II_{.3}:}~E=k_{1}b(x_{1}),~F=k_{2}b(x_{1})e^{-x_{2}},~G=b(x_{1})e^{-2x_{2}}.$ (2.8bb)
Rotation constants are
$\rho=-b_{,1}/(2b),~\sigma=\nu=0,~\tau=-i/(\sqrt{2b}\sqrt{k_{1}-k_{2}^{2}}).$ (2.8bc)
There exist no invariants. The surfaces $x_{1}=const$ have the negative Gaussian curvature $K=-1/(b(k_{1}-k_{2}^{2}))$.
If however $b_{,1}=0$, the Gaussian curvature $K$ is an invariant and the metric (2.8bb) degenerates to a horizon. (2.8bb) is then equivalent to (2.8ag).
6.4 Non-Abelian groups $G_{2}II$ with spacelike generators coplanar with the null direction
Finally, the case $\beta=0$ has to be treated. Here the changes
$\alpha^{\prime}=f_{1}\alpha,~\chi^{\prime}=\chi-h(x_{2})$ (2.8bd)
are possible, leading to (since obviously $\alpha\neq 0$) $\alpha^{\prime}=1,\chi^{\prime}=0$. The KE can be integrated with the result
$E=E_{0}(x_{2}),~F=F_{0}(x_{2})e^{-x_{1}},~G=G_{0}(x_{2})e^{-2x_{1}}.$ (2.8be)
The coordinate transformations
$x_{1}^{\prime}=x_{1}+f(x_{2}),~x_{2}^{\prime}=g(x_{2}),~x_{3}^{\prime}=x_{3},$ (2.8bf)
leave this metric structure invariant, but the functions $E_{0},F_{0},G_{0}$ change as
$E^{\prime}_{0}=E_{0}/g_{2}^{2},~F^{\prime}_{0}=F_{0}e^{f}/g_{2},~G^{\prime}_{0}=G_{0}e^{2f}.$ (2.8bg)
With suitable $f(x_{2}),g(x_{2})$ we can obtain a certain normal form ($k>1$):
$\mathbb{G_{3}VIII._{4}}:E=1,~F=e^{-x_{1}},~G=k(x_{2})e^{-2x_{1}}.$ (2.8bh)
An integration constant was absorbed in a changed $x_{2}$. The residual transformations leaving this normal form invariant are
$x^{\prime}_{1}=x_{1},~x^{\prime}_{2}=x_{2}+const,~x^{\prime}_{3}=x_{3}.$ (2.8bi)
The rotation coefficients for (2.8bh) are
$\rho=\frac{1}{2},~|\sigma|=\frac{1}{2}\frac{\sqrt{k}}{\sqrt{k-1}},~\sigma=\frac{1}{2}+\frac{i}{2\sqrt{k-1}},$ (2.8bj)
$\nu=-\frac{1}{2\sqrt{k-1}},~\tau=\frac{ik_{,2}}{2^{3/2}\sqrt{k(k-1)}}.$ (2.8bk)
The invariants are
$j=\sqrt{1-1/k},~I_{1}=-2/\sqrt{k},~I_{2}=I_{3}=J=0,$ (2.8bl)
$L=\frac{k_{,2}e^{is/2}}{(k-1)\sqrt{2k}},~M=-i\frac{k_{,2}e^{is/2}}{(k-1)^{3/2}\sqrt{2k}},$ (2.8bm)
note $e^{is/2}=(\sigma/|\sigma|)^{1/2}$. The surfaces $x_{1}=const$ have the Gaussian curvature
$K=-\frac{k_{,22}}{2(k-1)}+\frac{(k_{,2})^{2}}{4(k-1)^{2}}.$ (2.8bn)
We may ask if the metric (2.8bh) - as its notation suggests - admits an additional KV, say $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(A,B,C)$. The KE for this vector represents a system of differential equations for $A,B,C$, solved by
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(C_{,3},0,C),$ (2.8bo)
$C$ depends only on $x_{3}$. The commutator relations are given by
$|X_{1},X_{2}|=X_{1},~|X_{1},X_{3}|=(C_{,33},0,C_{,3}),~|X_{2},X_{3}|=(x_{3}C_{,33},0,x_{3}C_{,3}-C).$ (2.8bp)
If $C$ is a constant, $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$ essentially reduces to $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}$. The choice $C=x_{3}$ reproduces $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$, but $C=x_{3}^{2}$ or $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(2x_{3},0,x_{3}^{2})$ leads to the commutator relations
$|X_{1},X_{2}|=X_{1},~|X_{1},X_{3}|=2X_{2},~|X_{2},X_{3}|=X_{3},$ (2.8bq)
corresponding to a three-parameter Lie group of the Bianchi type VIII, with operators $X_{i}$ as given by Petrov [30].
7 Groups $G_{3}$ with spacelike generators
7.1 Overview of Groups $G_{3}$
The classification of three-parameter Lie groups was given by Bianchi [6], see also Petrov [30]:
$I.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=0,~[X_{2},X_{3}]=0$
$II.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=0,~[X_{2},X_{3}]=X_{1}$
$III.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=X_{1},~[X_{2},X_{3}]=0$
$IV.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=X_{1},~[X_{2},X_{3}]=X_{1}+X_{2}$
$V.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=X_{1},~[X_{2},X_{3}]=X_{2}$ (2.8a)
$VI.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=X_{1},~[X_{2},X_{3}]=qX_{2}~(q\neq 0,1)$
$VII.~[X_{1},X_{2}]$ $=$ $0,~[X_{1},X_{3}]=X_{2},~[X_{2},X_{3}]=-X_{1}+qX_{2}~(q^{2}<4)$
$VIII.~[X_{1},X_{2}]$ $=$ $X_{1},~[X_{1},X_{3}]=2X_{2},~[X_{2},X_{3}]=X_{3}$
$IX.~[X_{1},X_{2}]$ $=$ $X_{3},~[X_{1},X_{3}]=-X_{2},~[X_{2},X_{3}]=X_{1}$
The groups $I$ through $VII$ admit an Abelian $G_{2}$ as subgroup. We can use the results of sections 6.1 and 6.2. Thus we assume $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3}$ and $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=\delta^{i}_{2}$ as the first two commuting Killing fields not coplanar with the null direction (denoted as first cases). Then the resulting metric components $E,F,G$ depend only on a single coordinate $x_{1}$.
We have also to ask for every Bianchi class, if KVs coplanar with the null direction do exist (denoted as second cases). We then follow section 6.2 and write the first two KVs as $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3}$ and $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(1,0,\gamma(x_{2}))$.
The third Killing field is always written $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(A,B,C)$. The calculations will also show whether there are any null vectors among the Killing vectors, i.e. whether the corresponding metrics represent horizons.
7.2 $G_{3},$ Bianchi type I (first case)
Assume $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3}$ and $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=\delta^{i}_{2}$. The commutator relations $|X_{1},X_{3}|=0,~|X_{2},X_{3}|=0$ state that the components $A,B,C$ of the third KV can depend at most on $x_{1}$. The first part of the KE (2.8b) for the third KV shows that $B$ and $C$ are constants. From the second part (2.8c) one obtains
$AE_{,1}=0,~AF_{,1}=0,~AG_{,1}=0$ (2.8b)
i.e. a constant metric in case $A\neq 0$. $A=0$ is ruled out, since the third KV would be a linear combination of the first two: $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=B\mathrel{{\mathop{\xi}\limits_{2}}}^{i}+C\mathrel{{\mathop{\xi}\limits_{1}}}^{i}$, and no $G_{3}$ exists. With the residual transformation (2.8h) we obtain $A=1$. With $X^{\prime}_{3}=X_{3}-BX_{2}-CX_{1}$ we can also remove the components $B$ and $C$ from the third KV without violation of the commutator relations, thus $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=\delta^{i}_{1}$, which is null. We have the horizon (2.8ae) with a vanishing invariant $K$.
7.3 $G_{3},$ Bianchi type I (second case)
Are KVs possible for the Bianchi type I, which are coplanar with the null direction? Taking $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3},~\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=(1,0,\gamma(x_{2}))$, a third Killing field commuting with the first two requires that the components $A,B,C$ depend only on $x_{2}$ and that $B\gamma_{,2}=0$. The latter condition shows that three cases must be considered:
$\gamma_{,2}\neq 0,B=0,$
$\gamma_{2}=0,B\neq 0,$
$\gamma_{2}=B=0.$
Assuming first $\gamma_{,2}\neq 0$, $B=0$. $A\neq 0$ is required, since otherwise $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=C\mathrel{{\mathop{\xi}\limits_{1}}}^{i}$ and only a $G_{2}$ exists. With $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(A(x_{2}),0,C(x_{2}))$ we write the KE
$AE_{,1}+2FC_{,2}=0,~AF_{,1}+GC_{,2}=0,~AG_{,1}=0.$
Using the representation (6.2) of the metric, we find the further restriction $A\gamma_{,2}+C_{,2}=0.$ Since $\gamma_{,2}\neq 0$, we have $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(C_{,2}/\gamma_{,2},0,C).$ The third KV depends on a - by means of (6.2) - non-removable arbitrary function $C(x_{2})\neq\gamma$. This case was already treated, see (2.8v).
The next case is $\gamma_{,2}=0,B\neq 0$. Here a look at (6.2) shows that the components of the metric depend only on $x_{2}$. The KEs for $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$ are
$E_{,2}B+2EB_{,2}+2FC_{,2}=0,~F_{,2}B+FB_{,2}+GC_{,2},~G_{,2}B=0.$ (2.8c)
Since $B\neq 0$, we have $G=b=const$. The equations (2.8c) can be further integrated. With constants $a,b,c$ and functions $l(x_{2}),m(x_{2})$ we have
$E=l^{2}(a+bm^{2}-2cm),~F=l(c-bm),~G=b.$
However, the functions $l,m$ are spurious and can be removed by a coordinate transformation (6.2): $f_{,2}=-l(x_{2})A(x_{2}),~g_{,2}=l(x_{2}),~h_{,2}=-m(x_{2})l(x_{2})$. The final result is E=a,F=c,G=bE=a,~F=c,~G=b, that is a constant metric, and $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(0,1,0)$, corresponding to the horizon metric (2.8ae).
Finally we have the case $\gamma_{,2}=0,B=0$. Again the metric depends only on $x_{2}$:
$E=k(x_{2}),~F=l(x_{2}),~G=m(x_{2}).$ (2.8d)
The third KV is a combination of the first
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=A\mathrel{{\mathop{\xi}\limits_{2}}}^{i}+(C-\gamma)\mathrel{{\mathop{\xi}\limits_{1}}}^{i},$
where $C$ (like $\gamma$) is a constant from the KE for $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$, thus no $G_{3}$ with spacelike generators exists here. The metric admits however an isotropic KV, connected to $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}$ and $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ via $\epsilon^{i}=\gamma\mathrel{{\mathop{\xi}\limits_{1}}}^{i}-\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$. It is therefore also a horizon, identical to metric (2.8af) with slightly changed coordinates.
7.4 $G_{3}$, Bianchi type II
We first assume that $X_{1}$ and $X_{2}$ are not coplanar with the null direction, hence $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3}$ and $\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=\delta^{i}_{2}$. $|X_{1},X_{3}|=0$ says that $A,B,C$ are independent of $x_{3}$, and $|X_{2},X_{3}|=\delta^{i}_{3}$ gives the further restriction $(A,B,C)=(A_{0}(x_{1}),b_{0},x_{2}+c_{0})$ with constants $b_{0},c_{0}$. $A_{0}\neq 0$ is required for a non-singular metric. With the transformation (2.8h) we have $A_{0}=1,~c_{0}=0$, also $b_{0}=0$ can be assumed without loss of generality. Then the KE for $\mathrel{{\mathop{\xi}\limits_{3}}}$ leads to
$E_{,1}+2F=0,~F_{,1}=-G,~G_{,1}=0,$
and integrated to
$\mathbb{G_{3}II_{.1}:}$
$E$ $=$ $a-2cx_{1}+bx_{1}^{2},$
$F$ $=$ $c-bx_{1},$ (2.8e)
$G$ $=$ $b.$
$a,b,c$ are integration constants. With a suitable coordinate change $x_{1}^{\prime}=x_{1}+const$, $c$ disappears. The rotation cofficients are:
$\rho=0,~|\sigma|=|b|/(2\sqrt{ab-c^{2}}),$ (2.8f)
$\sigma=-ib/(2\sqrt{ab-c^{2}}),\nu=b/(2\sqrt{ab-c^{2}}),\tau=0.$ (2.8g)
Invariants are
$j=0,~I_{1}=2,N=0,$ (2.8h)
the surfaces $x_{1}=const$ are flat. Notice that $X_{1}$ and $X_{3}$ are coplanar with the null direction, which was not to be expected from the outset. Nevertheless, we have still to deal with the second case that $X_{1},X_{2}$ is coplanar with the null direction. Again $|X_{1},X_{3}|=0$ implies that $A,B,C$ are independent of $x_{3}$. From $|X_{2}X_{3}|=\delta^{i}_{3}$ we have $A_{1}=0,~B\gamma_{,2}=-1$, hence $B\neq 0,\gamma_{,2}\neq 0$. The transformatiom (6.2) allows A=0,B=1,C=0A=0,~B=1,~C=0, hence $\gamma_{,2}=-1$. The KE for $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$ gives in combination with (6.2)
$E=a+2cx_{1}+bx_{1}^{2},~F=c+bx_{1},~G=b.$
But this is essentially the same metric as (7.4): The roles of $X_{1}$ and $X_{2}$ are reversed in the two cases.
7.5 $G_{3}$, Bianchi types III,V,VI (first case)
The Bianchi types III,V,VIIII,~V,~VI have the commutator relations
$|X_{1},X_{2}|=0,~~|X_{1},X_{3}|=X_{1},~~|X_{2},X_{3}|=qX_{2}$ (2.8i)
with a parameter $q$ that distinguishes the different types:
$III:q=0,~V:q=1,~VI:q\neq 0,1.$ (2.8j)
In this section we will treat them together. The condition $|X_{1},X_{3}|=X_{1}$ provides $A_{,3}=0,~B_{,3}=0,~C_{,3}=1$. The remaining condition $|X_{2},X_{3}|=qX_{2}$ finally leads to
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(A(x_{1}),~B=qx_{2}+c_{1},~C=x_{3}+c_{2})$ (2.8k)
with two integration constants $c_{1},c_{2}$. The KE for $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$ would lead to $G=0$, if $A=0$, hence $A\neq 0$ and, after a coordinate transformation (2.8h), $A=1$. (2.8h) also removes the constant $c_{2}$; $c_{1}$ can be removed if $q\neq 0$. Considering also (2.8i), the equations for the metric are finally:
$E_{,1}+2Eq=0,\\ F_{,1}+F(1+q)=0,\\ G_{,1}+2G=0.$ (2.8l)
The integration gives
$E=ae^{-2qx_{1}},~F=ce^{-(q+1)x_{1}},~G=be^{-2x_{1}}$ (2.8m)
with constants $a,b,c,~ab-c^{2}>0$. We now specialize in the different Bianchi types.
Bianchi type III,q=0III,~q=0, first case
Metric:
$\mathbb{G_{3}III{.1}:}$
$E=a,~F=ce^{-x_{1}},~G=be^{-2x_{1}},$ (2.8n)
$|g_{AB}|=(ab-c^{2})\exp(-2x_{1})$ (2.8o)
Rotation coefficients:
$\rho=\frac{1}{2},~\nu=-\frac{c}{2\sqrt{ab-c^{2}}},~\sigma=\frac{1}{2}-i\nu,~\tau=0.$ (2.8p)
Invariants:
$j=\sqrt{(ab-c^{2})/(ab)},$ (2.8q)
$I_{1}=-2c/\sqrt{ab},~I_{2}=0,~I_{3}=I_{1}^{2}-4(1-j^{2})=0,~M=0.$ (2.8r)
The surfaces $x_{1}=const$ are flat.
Bianchi type V,q=1V,~q=1, first case
Metric:
$\mathbb{G_{3}V_{.1}:}$
$E=ae^{-2x_{1}},~F=ce^{-2x_{1}},~G=be^{-2x_{1}}.$ (2.8s)
$|g_{AB}|=(ab-c^{2})\exp(-4x_{1}).$ (2.8t)
Rotation coefficients:
$\rho=1,~\sigma=0,~\nu=0,~\tau=0.$ (2.8u)
No invariants, the surfaces $x_{1}=const$ are flat.
Bianchi type $VI,~q\neq 0,1$, first case
The metric is (2.8m), denoted $G_{3}VI_{.1}$.
Rotation coefficients:
$\rho=(q+1)/2,~\sigma=(1-q)(1+ic/\sqrt{ab-c^{2}})/2,$ (2.8v)
$\nu=c(q-1)/(2\sqrt{ab-c^{2}}),~\tau=0.$ (2.8w)
Invariants:
$j=\frac{q+1}{q-1}\sqrt{1-c^{2}/(ab)},~I_{1}=2c/\sqrt{ab},~I_{2}=0,~J=0$ (2.8x)
7.6 $G_{3}$, Bianchi types III,V,VI (second cases)
If $\mathrel{{\mathop{\xi}\limits_{1}}}^{i},\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ are coplanar with the null direction, the commutator relations for the third KV lead to
$A=a_{0}(x_{2})+qx_{1},~B=b_{0}(x_{2}),~C=x_{3}+c_{0}(x_{2}),~\gamma(q-1)+b_{0}(x_{2})\gamma_{,2}=0.$
For a regular metric, $b_{0}(x_{2})$ must be nonzero. This follows from the KE with $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$. The coordinate transformations (6.2) allow to obtain $a_{0}(x_{2})=0,~b_{0}(x_{2})=1,~c_{0}(x_{2})=0$. We now write the KE for the simplified third KV:
$qx_{1}E_{,1}+E_{,2}=0,~qx_{1}F_{,1}+F_{,2}+F=0,~qx_{1}G_{,1}+G_{,2}+2G=0$
and compare the result with (6.2). Beside
$\gamma=ke^{x_{2}},k=const$ (2.8y)
we finally obtain
$E$ $=$ $a-2kc(1-q)x_{1}e^{-qx_{2}}+k^{2}b(1-q)^{2}x_{1}^{2}e^{-2qx_{2}},$
$F$ $=$ $ce^{-x_{2}}-kb(1-q)x_{1}e^{-(1+q)x_{2}},$ (2.8z)
$G$ $=$ $be^{-2x_{2}}$
with
$EG-F^{2}=(ab-c^{2})e^{-2x_{2}}.$ (2.8aa)
We specialize again in the different Bianchi types.
Bianchi type III,q=0III,~q=0, second case
Metric:
$\mathbb{G_{3}III_{.2}:}$
$E=a-2kcx_{1}+bk^{2}x_{1}^{2},$ (2.8ab)
$F=(c-bkx_{1})e^{-x_{2}},$ (2.8ac)
$G=be^{-2x_{2}}.$ (2.8ad)
Rotation coefficients:
$\rho=0,~\nu=bk/(2\sqrt{ab-c^{2}}),~\sigma=-i\nu,~\tau=i\sqrt{b/(2(ab-c^{2}))}$ (2.8ae)
Invariants:
$j=0,~I_{1}=2sgn(k),~N=4\sqrt{b}(\cos(s/2)-\sin(s/2))/\sqrt{2(ab-c^{2})}.$ (2.8af)
The surfaces $x_{1}=const$ have the constant negative Gaussian curvature
$K=-b/(ab-c^{2}).$ (2.8ag)
For $k=0$ the metric (7.6) degenerates into a horizon and can be transformed into (2.8ag).
Bianchi type V,q=1V,~q=1, second case
The metric is given by
$E=a,~F=ce^{-x_{2}},~G=be^{-2x_{2}}.$ (2.8ah)
Rotation coefficients:
$\rho=0,~\sigma=0,~\nu=0,~\tau=-i\sqrt{b}/\sqrt{2(ab-c^{2})}$ (2.8ai)
Invariants:
$K=-b/(ab-c^{2})<0.$ (2.8aj)
We have a horizon with the negative constant curvature $K$, transformable into (2.8ag).
Bianchi type $VI,~q\neq 0,1$, second case
The metric is given by (2.8z), denoted $G_{3}VI_{.2}$.
Rotation coefficients:
$\rho=0,~\sigma=-i\nu,~\nu=bk(1-q)\exp(-qx_{2})/(2\sqrt{ab-c^{2}}),$ (2.8ak)
$~\tau=-i\sqrt{b/(2(ab-c^{2}))}$ (2.8al)
Invariants:
$j=0,~I_{1}=2sgn(k(1-q)),$ (2.8am)
$N=-4\sqrt{b}(\cos(s/2)-\sin(s/2))/\sqrt{2(ab-c^{2})}.$ (2.8an)
The surfaces $x_{1}=const$ have the constant negative Gaussian curvature $K=-b/(ab-c^{2})$.
7.7 $G_{3}$, Bianchi type IV (first case)
[…]
$\mathbb{G_{3}IV_{.1}:}$
$E=(a-2cx_{1}+bx_{1}^{2})e^{-2x_{1}},$ (2.8ao)
$F=(c-bx_{1})e^{-2x_{1}},$ (2.8ap)
$G=be^{-2x_{1}}.$ (2.8aq)
Rotation coefficients:
$\rho=1,~\sigma=ib/(2\sqrt{ab-c^{2}}),~\nu=b/(2\sqrt{ab-c^{2}}),~\tau=0.$ (2.8ar)
Invariants:
$j=2\sqrt{ab-c^{2}}/|b|,~I_{1}=2,~I_{2}=0,~I_{3}=16(ab-c^{2})/b^{2},J=0,M=0.$ (2.8as)
7.8 $G_{3}$, Bianchi type IV (second case)
If $\mathrel{{\mathop{\xi}\limits_{1}}}^{i},\mathrel{{\mathop{\xi}\limits_{2}}}^{i}$ are coplanar with the null direction, the commutator relation $|X_{1},X_{3}|=X_{1}$ returns the same result as in the previous section, $A_{2}=0,~B_{,3}=0,~C_{,3}=1$. The second relation $|X_{2},X_{3}|=X_{1}+X_{2}$ provides $A_{,1}=1,~B\gamma_{,2}=-1$, so we have
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(x_{1}+a_{0}(x_{2}),B(x_{2}),x_{3}+c_{0}(x_{2})).$
Transformations (6.2) bring $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}$ into the form
$\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(x_{1},1,x_{3}).$
The KE for this KV yields:
$x_{1}E_{,1}+E_{,2}=0,$
$x_{1}F_{,1}+F_{,2}+F=0,$
$x_{1}G_{,1}+G_{,2}+2G=0.$
Using the equations (6.2) with $\gamma=-x_{2}+const$ in the KE, one can finally conclude that
$\mathbb{G_{3}IV_{.2}:}$
$E=a+2cx_{1}e^{-x_{2}}+bx_{1}^{2}e^{-2x_{2}},$ (2.8ata)
$F=ce^{-x_{2}}+bx_{1}e^{-2x_{2}},$ (2.8atb)
$G=be^{-2x_{2}},$ (2.8atc)
$EG-F^{2}=(ab-c^{2})e^{-2x_{2}}.$ (2.8atd)
Rotation coefficients:
$\rho=0,~\sigma=-i\nu,~\nu=-b\exp(-x_{2})/(2\sqrt{ab-c^{2}}),$ (2.8ataua)
$\tau=-i\sqrt{b}/\sqrt{2(ab-c^{2})}$ (2.8ataub)
Invariants:
$j=0,~I_{1}=-2,~N=-4\sqrt{b}(\cos(s/2)-\sin(s/2))/\sqrt{2(ab-c^{2})}$ (2.8atauav)
The Gaussian curvature of the surfaces $x_{1}=const$ is the negative constant $K=-b/(ab-c^{2})$.
7.9 $G_{3}$, Bianchi type VII (first case)
The commutator relations are satisfied with
[…]
The KE for the third KV can be written
$AE_{,1}+2qE-2F=0,~AF_{,1}+Fq+E-2F=0,~AG_{,1}+2F=0.$ (2.8atauaw)
In the case $A(x_{1})=0$ we have from the third equation $F=0$ and from the first equation $qE=0$. Since $E\neq 0$ for a nonsingular metric, $q=0$ is required. Thus a first case of a $GVII$ metric is
$\mathbb{G_{3}VII_{.1}:}$
$E=f(x_{1}),~F=0,~G=f(x_{1})$ (2.8atauax)
with a single function $f(x_{1})$. Rotation coefficients:
$\rho=-f_{,1}/(2f),\sigma=0,~\nu=0,~\tau=0,$ (2.8atauay)
No invariants for $f_{,1}\neq 0$. The surfaces $x_{1}=const$ are flat, $K=0$. If $f_{,1}=0$, $K$ is invariant, we have the horizon (2.8ae).
If $A(x_{1})\neq 0$, $A=1$ can be reached with a coordinate transformation $x^{\prime}_{1}=f(x_{1})$. The differential equations (2.8atauaw) are then integrated with the integration constants $a_{1},a_{2},a_{3}$:
$\mathbb{G_{3}VII_{.2}:}$
$E$ $=$ $e^{-qx_{1}}\bigl(a_{1}+a_{2}\cos(px_{1})+a_{3}\sin(px_{1})\bigr)$ (2.8atauaza)
$F$ $=$ $e^{-qx_{1}}\bigl(a_{1}q+(a_{2}q+a_{3}p)\cos(px_{1})$
$+(-a_{2}p+a_{3}q)\sin(px_{1})\bigr)/2$
$G$ $=$ $e^{-qx_{1}}\bigl(2a_{1}+(a_{2}[2-p^{2}]+a_{3}pq)\cos(px_{1})$
$+(a_{3}[2-p^{2}]-a_{2}pq)\sin(px_{1})\bigr)/2$
with $p^{2}+q^{2}=4$. The determinant simplifies to
$|g_{AB}|=p^{2}\exp(-2qx_{1})(a_{1}^{2}-a_{2}^{2}-a_{3}^{2})/4.$ (2.8atauazba)
We also have, using the abbreviations $\lambda=a_{2}(2-q^{2})-a_{3}pq,~\mu=a_{3}(2-q^{2})+a_{2}pq$,
$\rho=\frac{q}{2},~|\sigma|=\frac{p\sqrt{a_{2}^{2}+a_{3}^{2}}}{2\sqrt{a_{1}^{2}-a_{2}^{2}-a_{3}^{2}}},~\tau=0,$ (2.8atauazbb)
$\sigma_{1}$ $=$ $\frac{p(\mu\cos{px_{1}}-\lambda\sin{px_{1}})}{2(2a_{1}-\lambda\cos{px_{1}}-\mu\sin{px_{1}})},$ (2.8atauazbca)
$\sigma_{2}$ $=$ $\frac{pa_{1}}{2\sqrt{a_{1}^{2}-a_{2}^{2}-a_{3}^{2}}}+\frac{p\sqrt{a_{1}^{2}-a_{2}^{2}-a_{3}^{2}}}{\lambda\cos(px_{1})+\mu\sin(px_{1})-2a_{1}},$ (2.8atauazbcb)
$\nu=-\frac{pa_{1}}{2\sqrt{a_{1}^{2}-a_{2}^{2}-a_{3}^{2}}}+~\frac{p\sqrt{a_{1}^{2}-a_{2}^{2}-a_{3}^{2}}}{2a_{1}-\lambda\cos{px_{1}}-\mu\sin{px_{1}}}.$ (2.8atauazbcbd)
where $\sigma=\sigma_{1}+i\sigma_{2}$.
The invariants are constant:
$j$ $=$ $\frac{q}{p}\sqrt{a_{1}^{2}/(a_{2}^{2}+a_{3}^{2})-1},$ (2.8atauazbcbea)
$I_{1}$ $=$ $-\frac{2a_{1}}{\sqrt{a_{2}^{2}+a_{3}^{2}}},$ (2.8atauazbcbeb)
$I_{2}$ $=$ $0,$ (2.8atauazbcbec)
$I_{3}$ $=$ $16(a_{1}^{2}/(a_{2}^{2}+a_{3}^{2})-1)/p^{2},$ (2.8atauazbcbed)
$M$ $=$ $0.$ (2.8atauazbcbee)
The surfaces $x_{1}=const$ are flat. In the special case $a_{2}=a_{3}=0$ no invariants exist.
7.10 $G_{3}$, Bianchi type VII (second case)
Consider now the case where the first two Killing fields are coplanar with the null direction. Here we have
[…]
satisfying $X_{1},X_{3}]=0$. The remaining commutator relations $X_{1},X_{2}]=X_{2},~[X_{2},X_{3}]=-X_{1}+qX_{2}$ involve the third KV $X_{3}$ and lead to
$A=x_{3}+(q-\gamma)x_{1}+a_{0}(x_{2}),~B=b_{0}(x_{2}),~C=\gamma x_{3}+c_{0}(x_{2})$
and to the condition
$\gamma^{2}+1-q\gamma-\gamma_{,2}b_{0}(x_{2})=0.$ (2.8atauazbcbebf)
The metric is given by (6.2). With the transformation 6.2 we reach $a_{0}(x_{2})=0.$ $b_{0}$ transforms as $b^{\prime}_{0}(x^{\prime}_{2})=g_{,2}b_{0}(x_{2})$. Thus either $b_{0}(x_{2})=0$ holds or $b_{0}(x_{2})=1$ can be reached. In the case $b_{0}=0$ one concludes from (2.8atauazbcbebf) that (for $\gamma$ to be real) $q^{2}=4$ is required, which is not possible, hence $b_{0}=1$. Now the relation (2.8atauazbcbebf) can be integrated:
$\gamma=\frac{q}{2}+\frac{p}{2}\tan(\frac{p}{2}x_{2}).$ (2.8atauazbcbebg)
[…]
$\mathbb{G_{3}VII_{.3}:}$
$E$ $=$ $a+b\frac{p^{4}x_{1}^{2}e^{-qx_{2}}}{16\cos^{2}(px_{2}/2)}-c\frac{p^{2}x_{1}e^{-qx_{2}/2}}{2\cos(px_{2}/2)}$ (2.8atauazbcbebha)
$F$ $=$ $ce^{-qx_{2}/2}\cos(px_{2}/2)-bp^{2}x_{1}e^{-qx_{2}}/4$ (2.8atauazbcbebhb)
$G$ $=$ $be^{-qx_{2}}\cos^{2}(px_{2}/2)$ (2.8atauazbcbebhc)
and
$EG-F^{2}=(ab-c^{2})\cos^{2}(px_{2}/2)e^{-qx_{2}}.$ (2.8atauazbcbebhbi)
$a,b,c$ are constants with $ab-c^{2}>0$. The rotation coefficients are
$\rho=0,~\sigma=-ib\frac{p^{2}e^{-qx_{2}/2}}{8\cos(px_{2}/2)\sqrt{ab-c^{2}}}$ (2.8atauazbcbebhbj)
$\nu=b\frac{p^{2}e^{-qx_{2}/2}}{8\sqrt{ab-c^{2}}\cos{px_{2}/2}},~\tau=-i\frac{\sqrt{b}(q+p\tan(px_{2}/2))}{2\sqrt{2}\sqrt{ab-c^{2}}}.$ (2.8atauazbcbebhbk)
The invariants are:
$j=0,~I_{1}=2,~N=-2\sqrt{2b}\gamma(\cos(s/2)-\sin(s/2))/\sqrt{ab-c^{2}}.$ (2.8atauazbcbebhbl)
The Gaussian curvature of the two-dimensional surfaces $x_{1}=const$ is
$K=\frac{b}{2(ab-c^{2})}(2-q^{2}-pq\tan(px_{2}/2)).$ (2.8atauazbcbebhbm)
7.11 $G_{3},$ Bianchi type VIII
The first two KV’s of a $G_{3}VIII$ satisfy $|X_{1},X_{2}|=X_{1}$, corresponding to a $G_{2}II$. We assume that $X_{1}$ and $X_{2}$ are non-coplanar to the null direction. The results of section (6.3) can be adopted: We have $\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3},~\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=\delta^{i}_{2}+x_{3}\delta^{i}_{3}$ and the corresponding metric (2.8ao). For the third KV the communicator relations $|X_{1},X_{3}|=2X_{2},~|X_{2},X_{3}|=X_{3}$ give after integration
$A=k(x_{1})e^{x_{2}},~B=2x_{3}+le^{x_{2}},~C=x_{3}^{2}+me^{2x_{2}},$ (2.8atauazbcbebhbn)
where $k=k(x_{1})$ depends on $x_{1}$, $l,m$ are constants. The KE for $X_{3}$ yields a system of ordinary differential equations for the functions $a,b,c$, which are present in (2.8ao):
$a_{,1}k+2al+4cm=0,$ (2.8atauazbcbebhbo)
$b_{,1}k-2bl+4c=0,$ (2.8atauazbcbebhbp)
$c_{,1}k+2a+2bm=0.$ (2.8atauazbcbebhbq)
The remaining freedom in the choice of coordinates
$x_{1}^{\prime}=f(x_{1}),~x_{2}^{\prime}=x_{2}+g_{0},~x_{3}^{\prime}=x_{3}+h_{0}e^{x_{2}}$ (2.8atauazbcbebhbr)
makes it possible to change $k,l,m$:
$k^{\prime}(x_{,1}^{\prime})=e^{-g_{0}}f_{,1}k(x_{,1})$ (2.8atauazbcbebhbs)
$l^{\prime}=e^{-g_{0}}(l-2h_{0}),$ (2.8atauazbcbebhbt)
$m^{\prime}=e^{-2g_{0}}(m+lh_{0}-h_{0}^{2}).$ (2.8atauazbcbebhbu)
We first assume $k(x_{1})\neq 0$. By choosing $f_{,1}=e^{g_{0}}/k(x_{1})$, we reach $k^{\prime}(x^{\prime}_{1})=1$. With $h_{0}=l/2$ we have $l^{\prime}=0$, while $m$ changes to $m^{\prime}=e^{-2g_{0}}(m+l^{2}/4)$. Depending on the sign of $m+l^{2}/4$, there are three cases to consider, $m^{\prime}=\pm 1,~0$. After all these transformations have been carried out, (2.8atauazbcbebhbo)-(2.8atauazbcbebhbq) obtains the simplified form
$a_{,1}+4cm=0,$
$b_{,1}+4c=0,$ (2.8atauazbcbebhbv)
$c_{,1}+2a+2bm=0,$
with $m=\pm 1,0.$ Depending on $m$, there are three solutions to the system (2.8atauazbcbebhbv), giving rise to three types of metrics:
$m=1$
$\mathbb{G_{3}VIII._{1}}:$
$E=n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}},$ (2.8atauazbcbebhbw)
$F=e^{-x_{2}}(n_{1}e^{-4x_{1}}-n_{2}e^{4x_{1}}),$ (2.8atauazbcbebhbx)
$G=e^{-2x_{2}}(-n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}}),$ (2.8atauazbcbebhby)
$EG-F^{2}=(4n_{1}n_{2}-n_{0}^{2})e^{-2x_{2}}$ (2.8atauazbcbebhbz)
$n_{0},n_{1},n_{2},k=\sqrt{4n_{1}n_{2}-n_{0}^{2}}$ are integration constants. The rotation constants are
$\rho=0,~|\sigma|=4\sqrt{n_{1}n_{2}}/k,$ (2.8atauazbcbebhca)
$\nu=2\frac{4n_{1}n_{2}-n_{0}n_{1}e^{-4x_{1}}-n_{0}n_{2}e^{4x_{1}}}{k(-n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}})},$ (2.8atauazbcbebhcb)
$\sigma=\frac{2(n_{1}e^{-4x_{1}}-n_{2}e^{4x_{1}})}{-n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}}}-i\nu,$ (2.8atauazbcbebhcc)
$\tau=-i\frac{\sqrt{-n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}}}}{\sqrt{2}k}.$ (2.8atauazbcbebhcd)
The invariants are constant:
$j=0,~I_{1}=-\frac{n_{0}}{\sqrt{n_{1}n_{2}}}.$ (2.8atauazbcbebhce)
The surfaces $x_{1}=const$ have the negative Gaussian curvature
$K=(n_{0}-n_{1}e^{-4x_{1}}-n_{2}e^{4x_{1}})/k^{2}.$ (2.8atauazbcbebhcf)
$m=0$
$\mathbb{G_{3}VIII._{2}}:$
$E=a_{0},$
$F=(c_{0}-2a_{0}x_{1})e^{-x_{2}},$ (2.8atauazbcbebhcg)
$G=(b_{0}-4c_{0}x_{1}+4a_{0}x_{1}^{2})e^{-2x_{2}}.$
$EG-F^{2}=d_{0}e^{-2x_{2}}.$ (2.8atauazbcbebhch)
$a_{0},b_{0},c_{0},d_{0}=a_{0}b_{0}-c_{0}^{2}$ are constants. The rotation coefficients are
$\rho=0,|\sigma|=a_{0}/\sqrt{d_{0}}$ (2.8atauazbcbebhci)
$\sigma=\frac{2(c_{0}-2a_{0}x_{1})}{b_{0}-4c_{0}x_{1}+4a_{0}x_{1}^{2}}-i\nu,$ (2.8atauazbcbebhcj)
$\nu=\frac{a_{0}b_{0}-2c_{0}^{2}+4a_{0}c_{0}x_{1}-4a_{0}^{2}x_{1}^{2}}{\sqrt{d_{0}}(b_{0}-4c_{0}x_{1}+4a_{0}^{2}x_{1}^{2}},$ (2.8atauazbcbebhck)
$\tau=-i\frac{\sqrt{b_{0}-4c_{0}x_{1}+4a_{0}x_{1}^{2}}}{\sqrt{2}\sqrt{d_{0}}}.$ (2.8atauazbcbebhcl)
The invariants are:
$j=0,~I_{1}=-2,N=-4\frac{\sqrt{4a_{0}x_{1}^{2}+b_{0}-4c_{0}x_{1}}}{\sqrt{2(a_{0}b_{0}-c_{0}^{2})}}(\cos(s/2)-\sin(s/2))$ (2.8atauazbcbebhcm)
The surfaces $x_{1}=const$ have the negative Gaussian curvature
$K=(-b_{0}+4c_{0}x_{1}-4a_{0}x_{1}^{2})/d_{0}.$ (2.8atauazbcbebhcn)
m=1m=-1
$\mathbb{G_{3}VIII._{3}}:$
$E=n_{0}+n_{1}\sin(4x_{1})-n_{2}\cos(4x_{1})$
$F=(n_{1}\cos(4x_{1})+n_{2}\sin(4x_{1}))e^{-x_{2}}$ (2.8atauazbcbebhco)
$G=(n_{0}-n_{1}\sin(4x_{1})+n_{2}\cos(4x_{1}))e^{2x_{2}}$
$EG-F^{2}=d_{0}e^{-2x_{2}}$ (2.8atauazbcbebhcp)
$n_{0},n_{1},n_{3},d_{0}=n_{0}^{2}-n_{1}^{2}-n_{2}^{2}$ are constants with $d_{0}>0$ and $n_{1}^{2}+n_{2}^{2}\neq 0$. The rotation coefficients are
$\rho=0,~|\sigma|=2\sqrt{n_{1}^{2}+n_{2}^{2}}/\sqrt{d_{0}},$ (2.8atauazbcbebhcq)
$\sigma=\frac{2(n_{1}\cos(4x_{1})+n_{2}\sin(4x_{1})}{n_{0}-n_{1}\sin(4x_{1})+n_{2}\cos(4x_{1})}-i\nu,$ (2.8atauazbcbebhcr)
$\nu=-\frac{2(n_{1}^{2}+n_{2}^{2}-n_{0}n_{1}\sin(4x_{1})+n_{0}n_{2}\cos(4x_{1}))}{\sqrt{d_{0}}(n_{0}-n_{1}\sin(4x_{1})+n_{2}\cos(4x_{1}))}$ (2.8atauazbcbebhcs)
$\tau=-i\sqrt{n_{0}-n_{1}\sin(4x_{1})+n_{2}\cos(4x_{1})}/\sqrt{2d_{0}}$ (2.8atauazbcbebhct)
The invariants are constant:
$j=0,~I_{1}=-2n_{0}/\sqrt{n_{1}^{2}+n_{2}^{2}}$ (2.8atauazbcbebhcu)
The Gaussian curvature of the surfaces $x_{1}=const$ is negative:
$K=-(n_{0}-n_{1}\sin(4x_{1})+n_{2}\cos(4x_{1}))/d_{0}.$ (2.8atauazbcbebhcv)
In the special case $n_{1}=n_{2}=0$ divergence and shear vanish. The metric degenerates to a horizon with a forth KV $\mathrel{{\mathop{\xi}\limits_{4}}}^{i}=\delta^{i}_{1}$:
$E=n_{0},~F=0.~G=n_{0}e^{-2x_{2}},$ (2.8atauazbcbebhcw)
transformable into (2.8ai) with $a^{2}=1/n_{0}$.
We still have to deal with the case where $X_{1}$ and $X_{2}$ are coplanar with the null direction. Following section 5.4, we have
$\mathrel{{\mathop{\xi}\limits_{1}}}^{i}=\delta^{i}_{3},~\mathrel{{\mathop{\xi}\limits_{2}}}^{i}=\delta^{i}_{1}+x_{3}\delta^{i}_{3}$ (2.8atauazbcbebhcx)
for the first two KV. The metric can be written in a reduced form (2.8bh). The commutator relation $X_{1},X_{3}]=2X_{2}$ restricts the third KV $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(A,B,C)$:
$A=2x_{3}+a(x_{1},x_{2}),~B=b(x_{2}),~C=x_{3}^{3}+c(x_{2}).$ (2.8atauazbcbebhcy)
The remaining relation $X_{2},X_{3}]=X_{3}$ gives $a=a_{0}(x_{2})e^{x_{1}},~b=c=0$. The KE for $X_{3}$ lead to $a_{0}=0$. Thus the metric (2.8bh) admits a third KV $\mathrel{{\mathop{\xi}\limits_{3}}}^{i}=(2x_{3},0,x_{3}^{2})$, as already noted above.
At the beginning of this section, we assumed that the function $k(x_{1})$ was not zero. The case $k(x_{1})=0$ still needs to be dealt with. From (2.8atauazbcbebhbo)-(2.8atauazbcbebhbq) we obtain l=2c/b,m=a/b<0l=2c/b,m=-a/b<0. The transformations (2.8atauazbcbebhbt),(2.8atauazbcbebhbu) lead with suitable $g_{0},h_{0}$ to l=0,m=1l=0,~m=-1, i.e. to a=b,c=0a=b,~c=0:
$\mathbb{G_{3}VIII._{5}}:$
$E=a(x_{1}),~F=0,~G=a(x_{1})e^{-2x_{2}}$ (2.8atauazbcbebhcz)
$EG-F^{2}=a^{2}e^{-2x_{2}}.$
Rotation coefficients are
$\rho=-a_{,1}/(2a),~\sigma=0,~\nu=0,~\tau=-1/\sqrt{2a},$ (2.8atauazbcbebhda)
and the surfaces $x_{1}=const$ have the negative Gaussian curvature K=1/a.K=-1/a.
7.12 $G_{3}$, Bianchi type IX
The communicator relations for Bianchi type IX are:
$|X_{1}X_{2}|=X_{3},~|X_{1}X_{3}|=-X_{2},~|X_{2}X_{3}|=X_{1}.$
The first two relations fix the $x_{3}$ dependence of the KV components:
$\alpha$ $=$ $a_{1}\cos(x_{3})+a_{2}\sin(x_{3}),$
$\beta$ $=$ $b_{1}\cos(x_{3})+b_{2}\sin(x_{3}),$
$\gamma$ $=$ $c_{1}\cos(x_{3})+c_{2}\sin(x_{3}),$
$A$ $=$ $-a_{1}\sin(x_{3})+a_{2}\cos(x_{3}),$
$B$ $=$ $-b_{1}\sin(x_{3})+b_{2}\cos(x_{3}),$
$C$ $=$ $-c_{1}\sin(x_{3})+c_{2}\cos(x_{3}).$
$a_{i},b_{i},c_{i}$ generally depend on $x_{1}$ and $x_{2}$, but the first part of the KE (2.8b) shows that $b_{i}$ and $c_{i}$ are independent of $x_{1}$. The final commutator relation $|X_{2}X_{3}|=X_{1}$ leads to
$a_{1}a_{2,1}+b_{1}a_{2,2}-a_{2}a_{1,1}-b_{2}a_{1,2}-a_{1}c_{1}-a_{2}c_{2}=0,$ (2.8atauazbcbebhdb)
$b_{1}b_{2,2}-b_{2}b_{1,2}-b_{1}c_{1}-b_{2}c_{2}=0,$ (2.8atauazbcbebhdc)
$b_{1}c_{2,2}-b_{2}c_{1,2}-1-c_{1}^{2}-c_{2}^{2}=0.$ (2.8atauazbcbebhdd)
We may simplify the last three equations with help of the coordinate transformations (2.8y). Under (2.8y) the $a_{i},b_{i},c_{i}$ transform as
$a^{\prime}_{1}$ $=$ $\sin(h)(f_{,1}a_{2}+f_{,2}b_{2})+\cos(h)(f_{,1}a_{1}+f_{,2}b_{1})$ (2.8atauazbcbebhde)
$a^{\prime}_{2}$ $=$ $\cos(h)(f_{,1}a_{2}+f_{,2}b_{2})-\sin(h)(f_{,1}a_{1}+f_{,2}b_{1})$ (2.8atauazbcbebhdf)
$b^{\prime}_{1}$ $=$ $g_{,2}(b_{2}\sin(h)+b_{1}\cos(h)),$ (2.8atauazbcbebhdg)
$b^{\prime}_{2}$ $=$ $g_{,2}(b_{2}\cos(h)-b_{1}\sin(h)),$ (2.8atauazbcbebhdh)
$c^{\prime}_{1}$ $=$ $\sin(h)(c_{2}+b_{2}h_{,2})+\cos(h)(c_{1}+b_{1}h_{,2})$ (2.8atauazbcbebhdi)
$c^{\prime}_{2}$ $=$ $-\sin(h)(c_{1}+b_{1}h_{,2})+\cos(h)(c_{2}+b_{2}h_{,2}).$ (2.8atauazbcbebhdj)
We first consider the transformation of $b_{1}$ and $b_{2}$. From (2.8atauazbcbebhdd) follows that $b_{1}$ and $b_{2}$ cannot vanish simultaneously. Let $b_{1}\neq 0$. We may reach $b^{\prime}_{1}=1,b^{\prime}_{2}=0$ by choosing $h(x_{2})$ and $g(x_{2})$ as
$\tan(h)=b_{2}/b_{1},~g_{,2}=1/\sqrt{b_{1}^{2}+b_{2}^{2}}.$
The obtained values $b_{1}=1,~b_{2}=0$ are not changed for residual transformations with $h=0,~g_{,2}=1$. One obtains from (2.8atauazbcbebhdc) $c_{1}=0$ and from (2.8atauazbcbebhdd) $c_{2,2}=1+c_{2}^{2}$. Integration gives $c_{2}=\tan(x_{2}+const)$ or $c_{2}=\tan(x_{2})$ after removing the integration constant. The still available transformations $x^{\prime}_{1}=f(x_{1},x_{2}),x^{\prime}_{2}=x_{2},x^{\prime}_{3}=x_{3}$ lead to
$a^{\prime}_{1}=f_{,1}+f_{,2},~a^{\prime}_{2}=f_{,1}a_{2}.$ (2.8atauazbcbebhdk)
By solving the partial differential equation $f_{,1}a_{1}+f_{,2}=0$ we obtain $a^{\prime}_{1}=0$. $a_{1}$ remains zero provided $f_{,2}=0$. Note that equation (2.8atauazbcbebhdb) is already satisfied if $a_{2}=0$. However, a vanishing of $a_{2}$ cannot be reached by coordinate transformations since $f_{,1}\neq 0$. Thus the case $a_{2}=0$ must be treated separately.
$a_{2}=0$
We solve the second part of the KE (2.8c) for the metric components $E,F,G$. This gives
$\mathbb{G_{3}IX_{.1}:}$
$E=m(x_{1}),~F=0,~G=m(x_{1})\cos(x_{2})^{2}.$ (2.8atauazbcbebhdl)
$m(x_{1})$ is an arbitrary function of $x_{1}$. The rotation coefficients are
$\rho$ $=$ $-m_{,1}/(2m),$ (2.8atauazbcbebhdm)
$\sigma$ $=$ $0,~\nu=0,$ (2.8atauazbcbebhdn)
$\tau$ $=$ $-i\tan(x_{2})/\sqrt{2m}.$ (2.8atauazbcbebhdo)
The two-dimensional wave surfaces $x_{1}=const$ are spheres with the Gaussian curvature $K=1/m(x_{1})$. The light cone of the FLRW cosmologies is an example for this type of null hypersurfaces. In the special case $m=const$ the metric becomes the horizon $G_{\infty}\times G_{3}IV$, (2.8ai).
$a_{2}\neq 0$
If $a_{,2}\neq 0$, equation 2.8atauazbcbebhdb with $a_{1}=0$ leads to $a_{2,2}-a_{2}\tan(x_{2})=0$ and after integration to $a_{2}=a_{20}/\cos(x_{2})$, where $a_{20}$ might depend on $x_{1}$. But the transformation $a^{\prime}_{2}=f_{,1}a_{2}$ finally allows to put $a_{2}=1/\cos(x_{2})$. The second part of the KE (2.8c) gives after integration
$\mathbb{G_{3}IX_{.2}:}$
$E=l+m\sin(2x_{1})+n\cos(2x_{1}),$
$F=\cos(x_{2})(n\sin(2x_{1})-m\cos(2x_{1})),$ (2.8atauazbcbebhdp)
$G=\cos(x_{2})^{2}(l-m\sin(2x_{1})-n\cos(2x_{1})),$
$EG-F^{2}=(l^{2}-m^{2}-n^{2})\cos^{2}(x_{2})$ (2.8atauazbcbebhdq)
$l,m,n$ are integration constants here. We also note some rotation coefficients:
$\rho$ $=$ $0,$ (2.8atauazbcbebhdr)
$|\sigma|$ $=$ $\sqrt{m^{2}+n^{2}}/\sqrt{l^{2}-m^{2}-n^{2}},$ (2.8atauazbcbebhds)
$\nu$ $=$ $\frac{n^{2}+m^{2}-l(n\cos(2x_{1})+m\sin(2x_{1}))}{\sqrt{l^{2}-m^{2}-n^{2}}(l-m\sin(2x_{1})-n\cos(x_{1}))},$ (2.8atauazbcbebhdt)
$\tau$ $=$ $-\frac{i\tan(x_{2})\sqrt{l-m\sin(x_{1})-n\cos(x_{1})}}{\sqrt{2}\sqrt{l^{2}-m^{2}-n^{2}}}$ (2.8atauazbcbebhdu)
The only second-order invariant is
$I_{1}=\frac{2l}{\sqrt{m^{2}+n^{2}}}.$ (2.8atauazbcbebhdv)
The two-dimensional wave surfaces $x_{1}=const$ have the Gaussian curvature:
$K=\frac{l-n\cos(2x_{1})-m\sin(2x_{1})}{l^{2}-m^{2}-n^{2}}.$ (2.8atauazbcbebhdw)
In the case $m=n=0$ also this metric becomes the horizon metric (2.8ai).
8 Groups $G_{4}$ with spacelike generators
The groups of motion of order 4 are also classified in Petrov’s book [30]. Every $G_{4}$ contains a subgroup $G_{3}$, so we can make use of previous results: Every metric admitting a $G_{3}$ must be checked to see whether it permits an additional fourth KV $X_{4}$. $X_{4}$ is written as $\mathrel{{\mathop{\xi}\limits_{4}}}^{i}=(U,~V,~W)$. The first part of the KE requires $V_{,1}=0,~W_{,1}=0$.
8.1 Group $G_{4I}$
The group operators satisfy the relations
$|X_{1},X_{2}|=0,~|X_{2},X_{3}|=X_{1},~|X_{3},X_{1}|=0.$ (2.8atauazbcbebha)
$|X_{1},X_{4}|=kX_{1},~|X_{2},X_{4}|=X_{2},~|X_{3},X_{4}|=(k-1)X_{3}.$ (2.8atauazbcbebhb)
The subgroup is $G_{3II}$, hence metric (7.4) applies. The commutation relations with $X_{4}$ lead to
$U=(k-1)x_{1}=const,~V=x_{2},~W=kx_{3}=const.$ (2.8atauazbcbebhc)
One KE for $X_{4}$ gives $bk=0$, thus $k=0$. Another KE then leads to $ab-c^{2}=0$, corresponding to a singular metric. No compatible metric exists.
8.2 Group $G_{4II}$
Commutator relations:
$|X_{1},X_{2}|=0,~|X_{2},X_{3}|=X_{1},~|X_{3},X_{1}|=0,$ (2.8atauazbcbebhd)
$|X_{1},X_{4}|=2X_{1},~|X_{2},X_{4}|=X_{2},~|X_{3},X_{4}|=X_{2}+X_{3}.$ (2.8atauazbcbebhe)
Again $G_{3II}$ is subgroup, also here the KE for $X_{4}$ is not compatible with the commutator relation $|X_{3},X_{4}|=X_{2}+X_{3}$. No compatible metric exists.
8.3 Group $G_{4III}$
Commutator relations:
$|X_{1},X_{2}|=0,~|X_{2},X_{3}|=X_{1},~|X_{3},X_{1}|=0,$ (2.8atauazbcbebhf)
$|X_{1},X_{4}|=qX_{1},~|X_{2},X_{4}|=X_{3},~|X_{3},X_{4}|=-X_{2}+qX_{3}.$ (2.8atauazbcbebhg)
Also here a $G_{3II}$ is subgroup, and the commutator relations for $X_{4}$ cannot be satisfied. No compatible metric exists.
8.4 Group $G_{4IV}$
The commutator relations:
$|X_{1},X_{2}|=0,~|X_{1},X_{3}|=X_{1},~|X_{2},X_{3}|=0,$ (2.8atauazbcbebhh)
$|X_{1},X_{4}|=0,~|X_{2},X_{4}|=X_{2},~|X_{3},X_{4}|=0$ (2.8atauazbcbebhi)
differ from those given by Petrov, since we have interchanged $X_{4}$ und $X_{3}$. Thus $G_{3III}$ becomes a subgroup of $G_{4IV}$. Two metrics, (7.5) and (7.6) belong to $G_{3III}$. The first becomes singular, if the KE for $X_{4}$ is taken into account. For the second metric (7.6) the commutator relations for $X_{4}$ give
$U=x_{1}+f(x_{3}),~V=const,~W=const~e^{x_{2}},~k(V+1)=0.$ (2.8atauazbcbebhj)
The KE for $X_{4}$ leads to $bV=0$ and $bW_{2}=0$, hence (since $b\neq 0$) to V=0,W=0V=0,~W=0 and $k=0$. The metric is reduced to
$E=a,~F=ce^{-x_{2}},~G=be^{-2x_{2}}$ (2.8atauazbcbebhk)
and represents a horizon, equivalent to (2.8ag). $X_{2}$ becomes a null vector. The fourth KV $X_{4}$ is also null and can be written $\mathrel{{\mathop{\xi}\limits_{4}}}^{i}=(x_{1},0,0)$.
8.5 Group $G_{4V}$
Commutator relations:
$|X_{1},X_{2}|=0,~|X_{2},X_{3}|=X_{1},~|X_{3},X_{1}|=0,$ (2.8atauazbcbebhl)
$|X_{1},X_{4}|=2X_{1},~|X_{2},X_{4}|=X_{2},~|X_{3},X_{4}|=X_{2}+X_{3}.$ (2.8atauazbcbebhm)
Subgroup is $G_{3V}$. Two metrics are admitted, (7.5) and (2.8ah). For the first metric the commutator relations for $X_{4}$ give
$U=const,~V=x_{3},~W=-x_{2}$ (2.8atauazbcbebhn)
and the KE reads
caU=0,ba+2cU=0,cbU=0.c-aU=0,~b-a+2cU=0,~c-bU=0. (2.8atauazbcbebho)
This leads to c=0,a=b,U=0c=0,~a=b,~U=0. the metric becomes
$\mathbb{G_{4V.1}:}$
$E=ae^{-2x_{1}},~F=0,~G=ae^{-2x_{1}}.$ (2.8atauazbcbebhp)
$EG-F^{2}=a^{2}e^{-4x_{1}}.$ (2.8atauazbcbebhq)
The rotation coefficients are
$\rho=1,~\sigma=0,~\nu=0,~\tau=0$ (2.8atauazbcbebhr)
There are no invariants, the surfaces $x_{1}=const$ are flat.
The other metric (2.8ah) is not compatible with the KE for $X_{4}$.
8.6 Group $G_{4VI}$
The commutator relations are
$|X_{i},X_{j}|=0,~|X_{i},X_{4}|=C_{i}^{k}X_{k},$ (2.8atauazbcbebhs)
$i,j=1,2,3;~k=1,2,3,4$ (2.8atauazbcbebht)
$G_{4}VI$ contains a $G_{3}I$ group as subgroup. Since all $G_{3}I$ metrics are horizons, this also holds for $G_{4}VI$ metrics.
8.7 Group $G_{4VII}$
The commutator relations are
$|X_{1},X_{2}|=X_{1},~|X_{1},X_{3}|=2X_{2},~|X_{2},X_{3}|=X_{3},$ (2.8atauazbcbebhu)
$|X_{i},X_{4}|=0.~i=1,2,3.$ (2.8atauazbcbebhv)
The group $G_{4VII}$ contains a subgroup $G_{3VIII}$. Five metrics satisfy the symmetries of $G_{3VIII}$, but only $G_{3VIII._{4}}$ admits a space-like KV $X_{4}=\partial_{2}$ that satisfies the commutator relations. The KE for $X_{4}$ require $k=const>1$:
$\mathbb{G_{4VII}:}$
$E=1,~F=e^{-x_{1}},~G=ke^{-2x_{1}}.$ (2.8atauazbcbebhw)
The rotation coefficients are
$\rho=\frac{1}{2},~\sigma=\frac{1}{2}+\frac{i}{2\sqrt{k-1}},~\nu=-\frac{1}{2\sqrt{k-1}},~\tau=0.$ (2.8atauazbcbebhx)
Invariants are
$j=\sqrt{1-1/k},~I_{1}=-\frac{2}{\sqrt{k}},I_{2}=0,~M=0.$ (2.8atauazbcbebhy)
The Gaussian curvature of the $x_{1}=const$ surfaces vanishes.
8.8 Group $G_{4VIII}$
The commutator relations are
$|X_{1},X_{2}|=X_{3},~|X_{1},X_{3}|=-X_{2},~|X_{2},X_{3}|=X_{1}.$ (2.8atauazbcbebhz)
$|X_{i},X_{4}|=0.~i=1,2,3.$ (2.8atauazbcbebhaa)
$G_{4VIII}$ has $G_{3IX}$ as a subgroup and the associated metrics are (2.8atauazbcbebhdl) and (2.8atauazbcbebhdq). For (2.8atauazbcbebhdl), the commutator relations yield $U=f(x_{1}),V=0,W=0$. The KE limits the metric by $m_{,1}=0$, thus, with constant $m$,
$E=m,~F=0,~G=m\cos(x_{2})^{2}.$ (2.8atauazbcbebhab)
This is the metric of the horizon (2.8ad). If we start from the second metric (2.8atauazbcbebhdq), commutator relations and KE lead to the same final metric.
9 Final comments
A group $G_{r}$ is called transitive, if any two points on the manifold can be reached by group orbits. A well-known theorem [15] states: A necessary and sufficient condition that a group $G_{r}$ be transitive is that $r\geq n$ and that the rank of the matrix $(\mathrel{{\mathop{\xi}\limits_{s}}}^{i})~(i=1...n,s=1...r)$ is $n$. Since $n=3$ in our case, this applies only to the groups $G_{3}$ and $G_{4}$, the groups $G_{1}$ and $G_{2}$ are intransitive. The rank condition identifies the following $G_{3}$ metrics as intransitive:
$\mathbb{G_{3}VII._{1}},q=0,$ (2.8atauazbcbebha)
$\mathbb{G_{3}VIII._{5}}$ (2.8atauazbcbebhb)
Both manifolds are covered by two-dimensional surfaces $x_{1}=const$, on which the metric is transitive.
The set of all points a given point $x_{i}$ is mapped by the group operators of a $r$-parameter group $G_{r}$ is called the orbit of $x_{i}$. The orbits of the metrics discussed here can be determined relatively easily by integration. But this topic belongs to global issues that are not covered here.
Appendix A:  Classification of null hypersurfaces in terms of invariants
Geometrical properties of null hypersurfaces can be expressed in terms of differential invariants, as described in [11, 12]. The differential invariants - up to second order - are found in section 3 as triad-invariant functions of the rotation coefficients $\rho$ (divergence), $\sigma$ (shear) and $\tau,~\nu,~\chi$ and $\varphi$. Equivalently, the invariants are also uniquely determined by the inner metric $\gamma_{ik}$. Written as functions of the rotation coefficients, there is considerable freedom in their representation, depending on the choice of the coefficients. In this paper we have used adapted coordinates with $\gamma_{1A}=0$. A possible representation of rotation coefficients as functions of the inner metric in adapted coordinates is given below, together with the differential invariants up to second order. A classification of null hypersurfaces along these lines is given in Table 1.
Table 1: Classification of null hypersurfaces in term of invariants.
shear divergence conditions class invariants
$|\sigma|\neq 0$ $\rho\neq 0$ $I_{2}\neq 0$ 11 $j,I_{1},I_{2},J$
$I_{2}=0,~I_{1}^{2}-4(1-j^{2})\neq 0$ $2a$ $j,I_{1},I_{2}=0,J,M$
$I_{2}=0,~I_{1}^{2}-4(1-j^{2})=0$ $2b$ $j,I_{1},I_{2}=0,J=0,L,M$
$\rho=0$ $|I_{1}|=2$ 33 $j=0,~|I_{1}|=2,~N$
$|I_{1}|\neq 2$ 44 $j=0,~I_{1}$
$|\sigma|=0$ $\rho\neq 0$ 55 noinvariantsno~invariants
$\rho=0$ 66 $K$
With $\Gamma=|\gamma_{AB}|$ and the usual abbreviations $E=\gamma_{22},~F=\gamma_{23},~G=\gamma_{33}$ we have:
$\rho=-\Gamma_{,1}/(4\Gamma),$ (2.8atauazbcbebhc)
$\sigma=\Gamma_{,1}/(4\Gamma)-G_{,1}/(2G)+\frac{i}{2\sqrt{\Gamma}}(F_{,1}-FG_{,1}/G),$ (2.8atauazbcbebhd)
$\nu=(G_{,1}F-GF_{,1})/(2G\sqrt{\Gamma}),$ (2.8atauazbcbebhe)
$\tau=\frac{i}{2\sqrt{2G}}(G_{,3}/G-\Gamma_{,3}/\Gamma)+\frac{i}{2\sqrt{2\Gamma G}}(G_{,2}+FG_{,3}/G-2F_{,3}),$ (2.8atauazbcbebhf)
$\chi=\varphi=0.$ (2.8atauazbcbebhg)
The differential operators are given by
$D=\partial_{1},~\delta=\frac{i\sqrt{G}}{\sqrt{2\Gamma}}\partial_{2}+\frac{\sqrt{\Gamma}-iF}{\sqrt{2\Gamma G}}\partial_{3}.$ (2.8atauazbcbebhh)
The differential invariants in Table 1 are defined in terms of rotation coefficients by (notice $e^{is}=\sigma/|\sigma|$ and the abbreviations $\zeta=2j-iI_{1}+I_{2},~\vartheta=2j-iI_{1}-I_{2}$)
$j=\rho/|\sigma|,$ (2.8atauazbcbebhi)
$I=I_{1}+iI_{2}=\frac{i}{|\sigma|}\Bigl(\frac{D\rho}{\rho}-\frac{D\sigma}{\sigma}\Bigr)+\frac{2\nu}{|\sigma|},$ (2.8atauazbcbebhj)
$J=e^{is/2}\Bigl[(\vartheta\bar{\vartheta}-4)\Bigl(\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}\Bigr)+(\zeta\bar{\vartheta}-4)\Bigl(\frac{\delta\bar{\sigma}}{\bar{\sigma}}-\frac{\delta\rho}{\rho}\Bigr)+2\bar{\vartheta}(\zeta-\vartheta)\bar{\tau}\Bigr]$
$\qquad+e^{-is/2}\Bigl[2(\zeta-\vartheta)\Bigl(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\rho}{\rho}+2\tau\Bigr)\Bigr],$ (2.8atauazbcbebhk)
$K=\delta\tau+\bar{\delta}\bar{\tau}-2\tau\bar{\tau},$ (2.8atauazbcbebhl)
$L=e^{is/2}\Bigl[\frac{\delta\sigma}{\sigma}-\frac{\delta\rho}{\rho}-2\bar{\tau}\Bigr]+e^{-is/2}(j+i\sqrt{1-j^{2}})\Bigl[\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}-\frac{\bar{\delta}\rho}{\rho}-2\tau\Bigr],$ (2.8atauazbcbebhm)
$M=-2e^{is/2}\delta j/j,$ (2.8atauazbcbebhn)
$N=e^{is/2}(\frac{\delta\sigma}{\sigma}-\frac{\delta\bar{\sigma}}{\bar{\sigma}}-4\bar{\tau})/(1+i)-ie^{-is/2}(\frac{\bar{\delta}\sigma}{\sigma}-\frac{\bar{\delta}\bar{\sigma}}{\bar{\sigma}}+4\tau)/(1+i).$ (2.8atauazbcbebho)
We stress again: $j,I,J,K,L,M,N$ are invariants only, if certain conditions are satisfied for the rotation coefficients. These conditions can be taken from Table 1. E.g., $N$ is an invariant if and only if $|\sigma|\neq 0,~\rho=0$ and $|I_{1}|=2$.
Appendix B:  Overview of metrics with Killing symmetries
Table 2: Group $G_{\infty}$ (horizon)
metric generators invariants class
$E=E(x_{2},x_{3})$ $X_{0}=f(x_{1})\partial_{1}$ Gaussian 6
$F=F(x_{2},x_{3})$ curvature $K$
$G=G(x_{2},x_{3})$
Table 3: Group $G_{\infty}\times G_{1}I$ (horizon)
metric generators invariants class
$E=1$ $X_{0}=f(x_{1})\partial_{1}$ $K=$ 6
$F=0$ $X_{1}=\partial_{3}$ $(-2mm_{,22}+m_{,2}^{2})/m^{2}$
$G=m(x_{2})$
Table 4: Group $G_{\infty}\times G_{3}VII_{0}$ (horizon)
metric generators invariants class
$E=a^{2}$ $X_{0}=f(x_{1})\partial_{1}$ $K=0$ 6
$F=0$ $X_{1}=\partial_{3},X_{2}=\partial_{2}$
$G=a^{2}$ $X_{3}=x_{3}\partial_{2}-x_{2}\partial_{3}$
$a=const$
Table 5: Group $G_{\infty}\times G_{3}VIII$ (horizon)
metric generators invariants class
$E=a^{2}/x_{2}^{2}$ $X_{0}=f(x_{1})\partial_{1}$ $K=-1/a^{2}$ 6
$F=0$ $X_{1}=\partial_{3},~X_{2}=x_{2}\partial_{2}+x_{3}\partial_{3}$
$G=a^{2}/x_{2}^{2}$ $X_{3}=2x_{2}x_{3}\partial_{2}+(x_{3}^{2}-x_{2}^{2})\partial_{3}$
$a=const$
Table 6: Group $G_{\infty}\times G_{3}IX$ (horizon)
metric generators invariants class
$E=a^{2}$ $X_{0}=f(x_{1})\partial_{1}$ $K=1/a^{2}$ 6
$F=0$ $X_{1}=\partial_{3},~X_{2}=\cos{x_{3}}\partial_{2}+sin{x_{3}}\tan{x_{2}}\partial_{3}$
$G=a^{2}\cos{x_{2}}^{2}$ $X_{3}=-\sin{x_{3}}\partial_{2}+\cos{x_{3}}\tan{x_{2}}\partial_{3}$
$a=const$
Table 7: Group $G_{1}$
metric generators invariants class
$E=E(x_{1},x_{2})$ $X_{1}=\partial_{3}$ $j,I,J$ 1
$F=F(x_{1},x_{2})$ $(seeApp.A)$ $(in$
$G=G(x_{1},x_{2})$ $general)$
Table 8: Abelian Group $G_{2}I.1$
(space-like generators $X_{1},X_{2}$, not coplanar with null direction)
metric generators invariants class
$E=a(x_{1})$ $X_{1}=\partial_{3}$ $j=-d_{,1}/\sqrt{p}$ 1
$F=c(x_{1})$ $X_{2}=\partial_{2}$ $I_{1}=8(d/p)^{3/2}(a_{,11}[b_{,1}c-bc_{,1}]$
$G=b(x_{1})$ $b_{,11}[c_{,1}a-a_{,1}c]+c_{,11}[a_{,1}b-b_{,1}a])$
$d=ab-c^{2}$ $I_{2}=-4dj_{,1}/d_{,1}$
$p=(a_{,1}b-ab_{,1})^{2}+$ $J=0$
$4(a_{,1}c-ac_{,1})(b_{,1}c-bc_{,1})\neq 0$
Table 9: Abelian Group $G_{2}I.2$
(space-like generators $X_{1},X_{2}$, not coplanar with null direction)
metric generators invariants class
$E=a(x_{1})$ $X_{1}=\partial_{3}$ 5
$F=c(x_{1})$ $X_{2}=\partial_{2}$
$G=b(x_{1})$
$p=(a_{,1}b-ab_{,1})^{2}+$
$4(a_{,1}c-ac_{,1})(b_{,1}c-bc_{,1})=0$
Table 10: Abelian Group $G_{2}I.3$
(space-like generators $X_{1},X_{2}$, coplanar with null direction)
metric generators invariants class
$E=1+x_{1}^{2}\gamma_{,2}^{2}m(x_{2})$ $X_{1}=\partial_{3}$ $j=0$ 3
$F=x_{1}\gamma_{,2}m(x_{2})$ $X_{2}=\partial_{1}+\gamma\partial_{3}$ $I_{1}=-2\gamma_{,2}/|\gamma_{,2}|$
$G=m(x_{2})$ $N=(\cos(s)-sin(s))m,_{2}/(\sqrt{2}m)$
$\gamma_{,2}\neq 0$
Table 11: Non-Abelian Group $G_{2}II.1$
(space-like generators $X_{1},X_{2}$, not coplanar with null direction)
metric generators invariants class
$E=a(x_{1})$ $X_{1}=\partial_{3}$ $j=-d_{,1}/\sqrt{p}$ 1
$F=c(x_{1})e^{-x_{2}}$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$ $I_{1}=8(d/p)^{3/2}(a_{,11}[b_{,1}c-bc_{,1}]$
$G=b(x_{1})e^{-2x_{2}}$ $b_{,11}[c_{,1}a-ca_{,1}]+c_{,11}[a_{,1}b-b_{,1}a])$
$d=ab-c^{2}$ $I_{2}=2\frac{d}{\sqrt{p}}(2\frac{d_{,11}}{d_{,1}}-\frac{p_{,1}}{p})$
$p=(a_{,1}b-ab_{,1})^{2}$
[…]
$d_{,1}\neq 0$
Table 12: Non-Abelian Group $G_{2}II.2$
(space-like generators $X_{1},X_{2}$, not coplanar with null direction)
metric generators invariants class
$E=a(x_{1})$ $X_{1}=\partial_{3}$ $j=0$ 4
$F=c(x_{1})e^{-x_{2}}$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$ $I_{1}=-sgn(c)\frac{2B}{\sqrt{d}A^{3/2}}$
$G=b(x_{1})e^{-2x_{2}}$ $A=(a_{,1}b+b_{,1}a)^{2}-4a_{,1}b_{,1}c^{2}$
$d=ab-c^{2},~d_{,1}=0$ $B=4dc^{2}(a_{,11}b_{,1}-b_{,11}a_{,1})$
$ba_{,1}^{2}(a_{,1}b^{2}+b_{,1}ab-4b_{,1}c^{2})$
$ab_{,1}^{2}(b_{,1}a^{2}+a_{,1}ab-4a_{,1}c^{2})$
Table 13: Non-Abelian Group $G_{2}II.3$
(space-like generators $X_{1},X_{2}$, not coplanar with null direction)
metric generators invariants class
$E=k_{1}b(x_{1})$ $X_{1}=\partial_{3}$ 5
$F=k_{2}b(x_{1})e^{-x_{2}}$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$
$G=b(x_{1})e^{-2x_{2}}$
$b_{,1}\neq 0,k_{1}>k_{2}^{2}$
Table 14: Bianchi type $G_{3}$II.1
metric generators invariants class
$E=a-2cx_{1}+bx_{1}^{2}$ $X_{1}=\partial_{3}$ $j=0$ 3
$F=c-bx_{1}$ $X_{2}=\partial_{2}$ $I_{1}=2$
$G=b$ $X_{3}=\partial_{1}+x_{2}\partial_{3}$ $N=0$
$a,b,c=const$
Table 15: Bianchi type $G_{3}$III.1
metric generators invariants class
$E=a$ $X_{1}=\partial_{3}$ $j=\sqrt{1-c^{2}/(ab)}$ 2b
$F=ce^{-x_{1}}$ $X_{2}=\partial_{2}$ $I_{1}=-2c/\sqrt{ab},~I_{2}=0,~I_{3}=0$
$G=be^{-2x_{1}}$ $X_{3}=\partial_{1}+x_{3}\partial_{3}$ J=0,L=0,M=0J=0,~L=0,~M=0
$a,b,c=const$
Table 16: Bianchi type $G_{3}$III.2
metric generators invariants class
$E=a-2ckx_{1}+bk^{2}x_{1}^{2}$ $X_{1}=\partial_{3}$ $j=0$ 3
$F=(c-bkx_{1})e^{-x_{2}}$ $X_{2}=\partial_{1}+ke^{x_{2}}\partial_{3}$ $I_{1}=2sgn(k)$
$G=be^{-2x_{2}}$ $X_{3}=\partial_{2}+x_{3}\partial_{3}$ $N=4\sqrt{b}(\cos(s/2)-\sin(s/2))/$
$a,b,c,k=const$ $\sqrt{2(ab-c^{2})}$
Table 17: Bianchi type $G_{3}$IV.1
metric generators invariants class
$E=(a-2cx_{1}+bx_{1}^{2})e^{-2x_{1}}$ $X_{1}=\partial_{3}$ $j=2\sqrt{ab-c^{2}}/b$ 2a
$F=(c-bx_{1})e^{-2x_{1}}$ $X_{2}=\partial_{2}$ $I_{1}=2,~I_{2}=0$~
$G=be^{-2x_{1}}$ $X_{3}=\partial_{1}+x_{2}\partial_{2}+(x_{2}+x_{3})\partial_{3}$ $I_{3}=16(ab-c^{2})/b^{2}$
$a,b,c=const$ J=0,M=0J=0,~M=0
Table 18: Bianchi type $G_{3}$IV.2
metric generators invariants class
$E=a+2cx_{1}e^{-x_{2}}+bx_{1}^{2}e^{-2x_{2}}$ $X_{1}=\partial_{3}$ $j=0$ 3
$F=ce^{-x_{2}}+bx_{1}e^{-2x_{2}}$ $X_{2}=\partial_{1}-x_{2}\partial_{3}$ $I_{1}=-2$
$G=be^{-2x_{2}}$ $X_{3}=x_{1}\partial_{1}+\partial_{2}$ $N=-4\sqrt{b}(\cos(s/2)-\sin(s/2))$
$a,b,c=const$ $x_{3}\partial_{3}$ $\sqrt{2(ab-c^{2})}$
Table 19: Bianchi type $G_{3}$V.1
metric generators invariants class
$E=ae^{-2x_{1}}$ $X_{1}=\partial_{3}$ 5
$F=ce^{-2x_{1}}$ $X_{2}=\partial_{2}$
$G=be^{-2x_{1}}$ $X_{3}=\partial_{1}+x_{2}\partial_{2}+x_{3}\partial_{3}$
$a,b,c=const$
Table 20: Bianchi type $G_{3}$VI.1
metric generators invariants class
$E=ae^{-2qx_{1}}$ $X_{1}=\partial_{3}$ $j=\frac{q+1}{q-1}\sqrt{1-c^{2}/(ab)}$ 1
$F=ce^{-(1+q)x_{1}}$ $X_{2}=\partial_{2}$ $I_{1}=2/\sqrt{ab},~I_{2}=0$
$G=be^{-2x_{1}}$ $X_{3}=\partial_{1}+qx_{2}\partial_{2}+x_{3}\partial_{3}$ $I_{3}=16q(ab-c^{2})/(ab(q-1)^{2}),~J=0$
$q\neq 0,1$
$a,b,c=const$
Table 21: Bianchi type $G_{3}$VI.2
metric generators invariants class
$E=a-2ck(1-q)x_{1}e^{-qx_{2}}+$ $X_{1}=\partial_{3}$ $j=0$ 3
$bk^{2}(1-q)^{2}x_{1}^{2}e^{-2qx_{2}}$
$F=ce^{-x_{2}}-bk(1-q)x_{1}e^{-(1+q)x_{2}}$ $X_{2}=\partial_{1}+ke^{(1-q)x_{2}}\partial_{3}$ $I_{1}=2\frac{k(1-q)}{|k(1-q)|}$
$G=be^{-2x_{2}}$ $X_{3}=qx_{1}\partial_{1}+\partial_{2}+x_{3}\partial_{3}$ $N=-4\sqrt{b}(\cos(s/2)$
$q\neq 0,1,~a,b,c=const$ $\sin(s/2))\sqrt{2(ab-c^{2})}$
Table 22: Bianchi type $G_{3}$VII.1(intransitive group)
metric generators invariants class
$E=f(x_{1})$ $X_{1}=\partial_{3}$ 5
$F=0$ $X_{2}=\partial_{2}$
$G=f(x_{1})$ $X_{3}=x_{3}\partial_{2}-x_{2}\partial_{3}$
$q=0,f_{,1}\neq 0$
Table 23: Bianchi type $G_{3}$VII.2
metric generators invariants class
$E=e^{-qx_{1}}\bigl(a_{1}+a_{2}\cos(px_{1})+a_{3}\sin(px_{1})\bigr)$ $X_{1}=\partial_{3}$ $j=$ 2a
$F=e^{-qx_{1}}\bigl(a_{1}q+(a_{2}q+a_{3}p)\cos(px_{1})$ $X_{2}=\partial_{2}$ $\frac{q}{p}\sqrt{a_{1}^{2}/(a_{2}^{2}+a_{3}^{2})-1}$
$(-a_{2}p+a_{3}q)\sin(px_{1})\bigr)/2$ $I_{1}=-2a_{1}/\sqrt{a_{2}^{2}+a_{3}^{2}}$
$G=e^{-qx_{1}}\bigl(2a_{1}+(a_{2}[2-p^{2}]+a_{3}pq)\cos(px_{1})$ $X_{3}=\partial_{1}$ $I_{2}=0$
$(a_{3}[2-p^{2}]-a_{2}pq)\sin(px_{1})\bigr)/2$ $(x_{3}+qx_{2})\partial_{2}$ $I_{3}=16(a_{1}^{2}/(a_{2}^{2}+a_{3}^{2})$
$p,q,a_{1},a_{2},a_{3}=const,~p^{2}+q^{2}=4$ $x_{2}\partial_{3}$ $1)/p^{2},~J=0,~M=0$
Table 24: Bianchi type $G_{3}$VII.3
metric generators invariants class
$E=a+b\frac{p^{4}x_{1}^{2}e^{-qx_{2}}}{16\cos^{2}(px_{2}/2)}-c\frac{p^{2}x_{1}e^{-qx_{2}/2}}{2\cos(px_{2}/2)}$ $X_{1}=\partial_{3}$ $j=0$ 3
$F=ce^{-qx_{2}/2}\cos(px_{2}/2)-bp^{2}x_{1}e^{-qx_{2}}/4$ $X_{2}=\partial_{1}+\gamma\partial_{3}$ $I_{1}=2,~N=-2\sqrt{2b}\gamma$
$G=be^{-qx_{2}}\cos^{2}(px_{2}/2)$ $X_{3}=(x_{3}+(q-\gamma))\partial_{1}$ $(\cos(s/2)-\sin(s/2))$
$\gamma=\frac{q}{2}+\frac{p}{2}\tan(\frac{p}{2}x_{2}),~p^{2}+q^{2}=4$ $\partial_{2}+x_{3}\gamma\partial_{3}$ $\sqrt{ab-c^{2}}$
Table 25: Bianchi type $G_{3}$VIII.1
metric generators invariants class
$E=n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}}$ $X_{1}=\partial_{3}$ $j=0$ 4
$F=e^{-x_{2}}(n_{1}e^{-4x_{1}}-n_{2}e^{4x_{1}})$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$ $I_{1}=-\frac{n_{0}}{\sqrt{n_{1}n_{2}}}$
$G=e^{-2x_{2}}(-n_{0}+n_{1}e^{-4x_{1}}+n_{2}e^{4x_{1}})$ $X_{3}=e^{x_{2}}\partial_{1}+2x_{3}\partial_{2}+(x_{3}^{2}+e^{2x_{2}})\partial_{3}$
$n_{0},n_{1},n_{2}=const$
$4n_{1}n_{2}-n_{0}^{2}>0$
Table 26: Bianchi type $G_{3}$VIII.2
metric generators invariants class
$E=a_{0}$ $X_{1}=\partial_{3}$ $j=0$ 3
$F=(c_{0}-2a_{0}x_{1})e^{-x_{2}}$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$ $I_{1}=-2,~N=$
$G=(b_{0}-4c_{0}x_{1}+4a_{0}x_{1}^{2})e^{-2x_{2}}$ $X_{3}=e^{x_{2}}\partial_{1}+$ $\frac{4}{\sqrt{2}\sqrt{a_{0}b_{0}-c_{0}^{2}}}(\cos(s/2)-\sin(s/2))$
$a_{0},b_{0},c_{0}=const$ $2x_{3}\partial_{2}+x_{3}^{2}\partial_{3}$ $\sqrt{4a_{0}x_{1}^{2}+b_{0}-4c_{0}x_{1}}$
$a_{0}b_{0}-c_{0}^{2}>0$
Table 27: Bianchi type $G_{3}$VIII.3
metric generators invariants class
$E=n_{0}+n_{1}\sin(4x_{1})-n_{2}\cos(4x_{1})$ $X_{1}=\partial_{3}$ $j=0$ 4
$F=e^{-x_{2}}(n_{1}\cos(4x_{1})+n_{2}\sin(4x_{1}))$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$ $I_{1}=$
$G=e^{-2x_{2}}(n_{0}-n_{1}\sin(4x_{1})+n_{2}\cos(4x_{1}))$ $X_{3}=e^{x_{2}}\partial_{1}+2x_{3}\partial_{2}$ $2n_{0}/\sqrt{n_{1}^{2}+n_{2}^{2}}$
$n_{0},n_{1},n_{2}=const,~n_{0}^{2}>n_{1}^{2}+n_{2}^{2}\neq 0$ $(x_{3}^{2}-e^{-2x_{2}})\partial_{3}$
Table 28: Bianchi type $G_{3}$VIII.4
metric generators invariants class
$E=1$ $X_{1}=\partial_{3}$ $j=\sqrt{1-1/k}$ 2b
$F=e^{-x_{1}}$ $X_{2}=\partial_{1}+x_{3}\partial_{3}$ $I_{1}=-2/\sqrt{k},~I_{2}=I_{3}=J=0$
$G=k(x_{2})e^{-2x_{1}}$ $X_{3}=2x_{3}\partial_{1}+x_{3}^{2}\partial_{3}$ $L=k_{,2}e^{is/2}/\sqrt{2k(k-1)^{2}}$
$M=-ik_{,2}e^{is/2}/\sqrt{2k(k-1)^{3}}$
Table 29: Bianchi type $G_{3}$VIII.5 (intransitive group)
metric generators invariants class
$E=a(x_{1})$ $X_{1}=\partial_{3}$ 5
$F=0$ $X_{2}=\partial_{2}+x_{3}\partial_{3}$
$G=a(x_{1})e^{-2x_{2}}$ $X_{3}=2x_{3}\partial_{2}+(x_{3}^{2}-e^{2x_{2}})\partial_{3}$
Table 30: Bianchi type $G_{3}$IX.1
metric generators invariants class
$E=m(x_{1})$ $X_{1}=\partial_{3}$ 5
$F=0$ $X_{2}=\cos(x_{3})\partial_{2}+\sin(x_{3})\tan(x_{2})\partial_{3}$
$G=m(x_{1})\cos^{2}(x_{2})$ $X_{3}=-\sin(x_{3})\partial_{2}+\cos(x_{3})\tan(x_{2})\partial_{3}$
Table 31: Bianchi type $G_{3}$IX.2
metric generators invariants class
$E=l+m\sin(2x_{1})+n\cos(2x_{1})$ $X_{1}=\partial_{3}$ $j=0$ 4
$F=\cos(x_{2})(n\sin(2x_{1})$- $X_{2}=\frac{\sin(x_{3})}{\cos(x_{2})}\partial_{1}+$
$m\cos(2x_{1}))$ $\cos(x_{3})\partial_{2}+\sin(x_{3})\tan(x_{2})\partial_{3}$ $I_{1}=$
$G=\cos^{2}(x_{2})(l-m\sin(2x_{1})$- $X_{3}=\frac{\cos(x_{3})}{\cos(x_{2})}\partial_{1}$- $\frac{2l}{\sqrt{m^{2}+n^{2}}}$
$n\cos(2x_{1}))$ $\sin(x_{3})\partial_{2}+\cos(x_{3})\tan(x_{2})\partial_{3}$
$l,m,n=const$
Table 32: $G_{4}$V
metric generators invariants class
$E=ae^{-2x_{1}}$ $X_{1}=\partial_{3}$ 5
$F=0$ $X_{2}=\partial_{2}$
$G=ae^{-2x_{1}}$ $X_{3}=\partial_{1}+x_{2}\partial_{2}+x_{3}\partial_{3}$
$a=const$ $X_{4}=x_{3}\partial_{2}-x_{2}\partial_{3}$
Table 33: $G_{4}$VII
metric generators invariants class
$E=1$ $X_{1}=\partial_{3}$ $j=\sqrt{1-1/k}$ 2b
$F=e^{-x_{1}}$ $X_{2}=\partial_{1}+x_{3}\partial_{3}$ $I_{1}=-\frac{2}{\sqrt{k}},~I_{2}=0,~I_{3}=0$
$G=ke^{-2x_{1}}$ $X_{3}=2x_{3}\partial_{1}+x_{3}^{2}\partial_{3}$ $J=L=M=0$
$k=const>1$ $X_{4}=\partial_{2}$